### v1

Abstract
We present simple to implement Wald-type statistics that deliver a general nonparametric inference theory for linear restrictions on varying coefficients in a range of spatial autoregressive models. Our theory covers error dependence of a general form, allows for a degree of misspecification robustness via nonparametric spatial weights and permits inference on both varying regression and spatial coefficients. One application of our method finds evidence for constant returns to scale in the production function of the Chinese nonmetal mineral industry, while another finds a nonlinear impact of the distance to the employment center on housing prices in Boston. A simulation study confirms that our tests perform well in finite-samples.
JEL codes: C21, C31
Keywords: Spatial autoregression, varying coefficients, inference
1.   Introduction
This paper develops a general framework for inference on varying coefficient spatial autoregressive (SAR) models that is as easy to implement in practice as familiar Wald tests. SAR models are fast becoming a workhorse for researchers working in settings where economic agents interact via economic or social links. Such links can be geographic proximity, or proximity in some more general sense such as a social network. The SAR model can accommodate these types of general links easily, and this goes some way to explain its appeal. No surprise then that the baseline SAR model, due to Cliff and Ord (1981), has become much studied by econometricians. This model has long been popular with regional scientists, see Anselin (1988) for an early book length treatment; important econometric contributions that established the foundation for rigorous econometric analysis of such models include Kelejian and Prucha (1998) and Lee (2004).
In keeping with the growing importance of SAR models, various strands of the literature have explored how to widen their scope. One such is the focus on varying coefficient SAR models, wherein model parameters are allowed to be unknown functions of some observed economic variable(s). This permits a flexible effect of the regressors and spatial dependence on the outcome variable, and indeed varying coefficient models are abundant in statistics and econometrics (see e.g. Fan and Zhang, 1999, 2008). In the SAR context, varying coefficient models have been studied by Sun et al. (2014) and a sequence of important papers by Sun (2016), Malikov and Sun (2017), Sun and Malikov (2018) and Sun (2024), for example. More generally, semiparametric SAR models have attracted much interest, see e.g. Su and Jin (2010), Su (2012), Robinson and Thawornkaiwong (2012) and Zhang (2013), but their focus tends to be on inference on the parametric component.
In this paper, we show how inference on varying coefficients in a range of SAR models can be conducted just like familiar parametric Wald tests for linear restrictions. The key idea is that if series approximations are used for the nonparametric varying coefficients then the series coefficients can be employed for very easy practical inference. This is motivated by the scientific aim of simplification of nonparametric/semiparametric inference. Indeed, a growing literature stresses such ideas, see e.g Ackerberg et al. (2012), Gupta (2018) and Korolev (2019), amongst others.
[…]
Using this series expansion approach, we theoretically justify Wald-type test statistics for inference on varying coefficients. We cover a baseline high-order SAR model with varying regression coefficients but constant spatial lag parameters and extend the model to allow for spatial error dependence. The baseline model is similar to the one estimated by Sun et al. (2014), but also features a growing number of spatial lags and covariates. We then show how to allow for a degree of misspecification robustness by incorporating nonparametric spatial weights à la Pinkse et al. (2002). Our approach is also shown to work for models with varying coefficients on the spatial lags, in the spirit of Malikov and Sun (2017).
We then apply our methods on two different studies where varying coefficients have been shown to play an important role. In the first application, we test for constant returns to scale (CRS) in the production function of the Chinese nonmetal mineral industry. This can be seen as a follow-up application to that of Li et al. (2002), who sugges that their estimate of the returns to scale technology, which varies with managerial expense, is close to being constant (i.e. output elasticities of capital and labor sum to unity) for most observed managerial expense levels. However, they do not have a test for CRS and do not model spatial dependence. Our analysis confirms their observations as we cannot reject the CRS hypothesis using their model and in much richer settings. This suggests that CRS technology may be a salient feature of this industry.
Our second study is an illustration on the famous Boston housing data. We estimate a model similar to Sun et al. (2014), who used a model selection procedure to recommend that location should have a varying effect on house prices with respect to some other features of the property. We formally test for the varying location effects of these features using our statistics and indeed find them to be statistically significant.
In a set of Monte Carlo simulation studies, we show that our tests have excellent finite-sample performance. In addition to tests that are based on the asymptotic normal approximation, we also propose a bootstrap version of our tests; the bootstrap tests perform particularly well for small sample sizes. We experiment with a range of SAR models and spatial dependence structures to demonstrate the full range of applicability and generality of our approach.
The next section introduces our method in the context of a baseline higher-order SAR model. Section 3 extends the model to allow for spatial error dependence of Kelejian and Prucha (2007) type, while Section 4 allows for nonparametric spatial weights (Pinkse et al., 2002). In Section 5 we allow the coefficient on the spatial lag to also vary (Malikov and Sun, 2017) and show how our approach can be adapted to this setting. Section 6 presents our empirical applications to the Chinese nonmetal mineral industry and Boston house prices, while Section 7 contains the simulation study. Proofs are in the appendix, and a supplementary online appendix contains a number of additional simulation results.
2.   Baseline spatial autoregression
Consider a vector of unknown functions $\delta\left(\cdot\right)=\left(\delta_{1}\left(\cdot\right),\ldots,\delta_{d_{\delta}}\left(\cdot\right)\right)^{\prime}$, and the model
$y_{in}=\sum_{j=1}^{d_{\lambda}}\lambda_{j}w_{in,j}^{\prime}y_{n}+x_{in}^{\prime}\beta+p_{in}^{\prime}\delta\left(z_{in}\right)+\epsilon_{in},i=1,\ldots,n,$ (2.1)
where $y_{n}$ is the $n\times 1$ vector with typical element $y_{in}$, $w_{in,j}^{\prime}$ is the $i$th row of the spatial weight matrix $W_{jn}$, $j=1,\ldots,d_{\lambda}$, $d_{\lambda}\rightarrow\infty$ as $n\rightarrow\infty$, and $\epsilon_{in}$ is $i.i.d.$ with mean 0 and unit variance. Also, $\left(x_{in},p_{in},z_{in}\right)\in\mathcal{X}\times\mathcal{P}\times\mathcal{Z}\subseteq\mathbb{R}^{d_{\beta}\times d_{\delta}\times d_{z}}$, and $d_{\beta}\rightarrow 0$ as $n\rightarrow\infty$ but $d_{z}$ and $d_{\delta}$ are fixed. For all $i=1,\ldots,n$, $z_{in}$ is throughout the paper uncorrelated with $\epsilon_{in}$ while $x_{in}$ is allowed to be correlated with $\epsilon_{in}$.
Permitting $d_{\lambda}$ and $d_{\beta}$ to diverge with sample size allows a flexible modelling approach, with the model becoming richer as more observations become available to the practitioner. This approach has been studied in statistics since at least the work of Huber (1973), but in spirit the sequence of experiments considered by Le Cam (1960) provides an even earlier reference. Econometricians have also studied such models in least squares and GMM settings, see e.g. Andrews (1985), Koenker and Machado (1999) and Cattaneo et al. (2018), to name just three references. In a SAR setting, Gupta and Robinson (2015), Gupta (2018) and Gupta and Qu (2024) emphasize the advantages of this approach, especially when clustered data naturally imply asymptotic regimes in which $d_{\lambda}$ diverges with $n$. It is worth noting that the results stated in this paper also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed but we state our theorems for the more complex case where they diverge.
We now mostly drop $n$ subscripting, but occasionally remind the reader of the dependence on sample size of various quantities. The paper takes the spatial weight matrices $W_{j}$ to be non-stochastic everywhere except in Section 4. In all sections these weight matrices are also uniformly bounded in row and column sums, a commonly used restriction to control spatial dependence (Kelejian and Prucha, 1998; Lee, 2004). This property is assumed to hold almost surely in Section 4, where the spatial weights are stochastic.
Our aim is to do inference on $\delta(\cdot)$, specifically we wish to test a fixed number $m$ restrictions of the form $S\delta(z)=s,z\in\mathcal{Z}$, where $S$ is a known, constant $m\times d_{\delta}$ matrix and $s$ is a known, constant $m\times 1$ vector. Because we can convert any linear restriction to an exclusion restriction, we will focus on tests of the null hypothesis
$H_{0}^{true}:\delta(z)=0,z\in\mathcal{Z}.$ (2.2)
Our test statistics will approximate $\delta(\cdot)$ by a series expansion and test if the coefficients in the expansion are jointly zero, thus converting the restriction in $H_{0}^{true}$ to an increasing-dimensional one, up to some suitably negligible approximation error.
2.1   Test statistic
We approximate $\delta_{k}(z)$ by $\psi_{k}^{h_{k}}(z)^{\prime}\alpha_{k}^{h_{k}}$, where $\psi_{k}^{h_{k}}(z)=\left(\psi_{k1}(z),\ldots,\psi_{kh_{k}}(z)\right)^{\prime}$ and $\alpha_{k}^{h_{k}}=\left(\alpha_{k1},\ldots,\alpha_{kh_{k}}\right)^{\prime}$, for some basis functions $\psi_{k\ell}(\cdot)$, $\ell=1,\ldots,h_{k}$, $k=1,\ldots,d_{\delta}$. Thus each $\delta_{k}(z)$ is approximated by a linear combination of $h_{k}$ basis functions $\psi_{k}^{h_{k}}$ with coefficients $\alpha_{k}^{h_{k}}$. Define the $d_{\alpha}\times 1$ vector $\psi_{i}\equiv\psi_{i}\left(p_{i},z_{i}\right)=\left(p_{i1}\psi_{1}^{h_{1}}\left(z_{i}\right)^{\prime},\ldots,p_{id_{\delta}}\psi_{d_{\delta}}^{h_{d_{\delta}}}\left(z_{i}\right)^{\prime}\right)^{\prime}$ and $\alpha=\left(\alpha_{1}^{h_{1}}{{}^{\prime}},\ldots,\alpha_{d_{\delta}}^{h_{d_{\delta}}}{{}^{\prime}}\right)^{\prime}$, with $d_{\alpha}=\sum_{k=1}^{d_{\delta}}h_{k}$. Then we can write (2.1) as
$y_{i}=\sum_{j=1}^{d_{\lambda}}\lambda_{j}w_{i,j}^{\prime}y+x_{i}^{\prime}\beta+\psi_{i}^{\prime}\alpha+u_{i},i=1,\ldots,n,$ (2.3)
with $u_{i}=r_{i}+\epsilon_{i}$, where $r_{i}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$ is the approximation error. Let $X$ and $\Psi$ be matrices with typical rows $x_{i}^{\prime}$ and $\psi_{i}^{\prime}$, and $u$ be with elements $u_{i}$. Writing $Q$ for the $n\times d_{\lambda}$ matrix with typical columns $W_{j}y$, (2.3) can be written in matrix notation as
$y=Q\lambda+X\beta+\Psi\alpha+u=L\xi+u,$ (2.4)
where $L=[Q,X,\Psi]$ and $\xi=\left(\lambda^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$.
The null hypothesis to approximate $H_{0}^{true}$ is
$H_{0}:\alpha=0,$ (2.5)
and our test will be implemented using the 2-stage least squares (2SLS) estimator
$\hat{\xi}\equiv\left(\hat{\lambda}^{\prime},\hat{\beta}^{\prime},\hat{\alpha}^{\prime}\right)^{\prime}=\left(L^{\prime}{P_{K_{1}}}L\right)^{-1}L^{\prime}{P_{K_{1}}}y,$ (2.6)
where $P_{K_{1}}=K_{1}\left(K_{1}^{\prime}K_{1}\right)^{-1}K_{1}^{\prime}$, with $K_{1}$ being an $n\times J_{1}$ instrument matrix with $J_{1}\geq d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$ but having the same asymptotic order i.e. ${J_{1}}/{d_{\xi}}$ tends to a constant that is at least unity, as $n\rightarrow\infty$. Let $T_{1}=\frac{1}{n}L^{\prime}{P_{K_{1}}}L$ and $\mathscr{D}_{1}=\frac{1}{n}RT_{1}^{-1}R^{\prime}$, where $R=\left[0_{d_{\alpha}\times\left(d_{\lambda}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $I_{d_{\alpha}}$ is the $d_{\alpha}\times d_{\alpha}$ identity matrix. . Then, the test statistic is
$\mathbb{W}_{1}=\frac{n\hat{\alpha}^{\prime}\mathscr{D}_{1}^{-1}\hat{\alpha}-d_{\alpha}}{\sqrt{2d_{\alpha}}}.$ (2.7)
2.2   Asymptotic properties
We begin with restrictions on the moments of various objects, noting that throughout the paper $C$ denotes a generic positive constant, arbitrarily large but independent of $n$. Let $k_{ri}$ be the elements of the instrumental variable matrix, which corresponds to $K_{1},K_{2}$ and $K_{3}$ as defined in this section, 4 and 5, respectively, and $l_{ri}$ denote the elements of $L$.
Assumption 1.
$\sup_{i\geq 1}\mathrm{E}\left(r_{i}^{2}\right)=o(n^{-1})$.
Assumption 2.
$\mathrm{E}\left|\epsilon_{i}\right|^{q}<C$, for some $q>4$.
Assumption 3.
$\mathrm{E}\left(l_{ri}^{2}\right)<C$ and $\mathrm{E}\left(k_{ri}^{2}\right)<C$.
Assumption 1 is a general assumption that controls the approximation error. Sufficient conditions for various cases can be found in Chen (2007). We avoid using specific rates for the approximation error because many types of approximation errors occur in the paper and explicitly allowing different rates of decay introduces extra notation without adding much insight beyond our general conditions. Assumption 2 is fairly standard when establishing a central limit theorem for quadratic forms and is used to check a Lyapunov condition. Assumption 3 imposes moment conditions on regressors and instruments.
To ease notation, we introduce the following definition which captures boundedness and non-multicollinearity properties of random or fixed matrices of growing dimension.
Definition 1.
For any square, symmetric and positive semi-definite matrix $A$, let $\overline{\alpha}(A)$ and $\underline{\alpha}(A)$ denote its largest and smallest eigenvalues, respectively. If $A$ is random, we say that $A$ has Property G if
$\overline{\alpha}(A)=O_{p}(1)\text{ and }\left\{\underline{\alpha}(A)\right\}^{-1}=O_{p}(1).$
If $A$ is non-random, we say that $A$ has Property G if
$\limsup_{n\rightarrow\infty}\overline{\alpha}(A)<\infty\text{ and }\liminf_{n\rightarrow\infty}\underline{\alpha}(A)>0.$
Assumption 4.
$n^{-1}K_{1}^{\prime}K_{1}$, $n^{-1}L^{\prime}K_{1}K_{1}^{\prime}L$ and $n^{-1}\Psi^{\prime}\Psi$ have 1.
We first show that the statistic $\mathbb{W}_{1}$ can be approximated by a quadratic form in $\epsilon$. Define $\mathcal{M}_{1}=\frac{1}{n}K_{1}\mathcal{V}_{1}K_{1}^{\prime}$, with $\mathcal{V}_{1}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{1}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$. Then, we have the following theorem.
Theorem 2.1.
Under $H_{0}$, Assumptions 1-4 and
$\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\xi}}{n}\rightarrow 0,$ (2.8)
as $n\rightarrow\infty$,
$\mathbb{W}_{1}-\frac{\epsilon^{\prime}\mathcal{M}_{1}\epsilon-d_{\alpha}}{\sqrt{2d_{\alpha}}}=o_{p}(1).$
We take this opportunity to remind the reader that the results stated in this paper also hold when $d_{\lambda}$ and $d_{\beta}$ are held fixed and we do not state separate theorems for that case.
Assumption 5.
$\left(I_{n}-\sum_{j=1}^{d_{\lambda}}\lambda_{j}W_{j}\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$.
Define $H_{\ell 1}\equiv H_{\ell 1,n}:\alpha=\alpha^{*}\equiv\nu_{1n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{\frac{1}{2}}$, with $\nu_{1n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{1n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in the proof of the next theorem. This sequence of local alternatives features a $d_{\alpha}^{1/4}$ damping factor that accounts for the cost of our nonparametric approach, and has been found in similar problems by Hong and White (1995) and Gupta (2018), amongst others. We can now state the main theorem of the section.
Theorem 2.2.
Under Assumptions 1-5 and
$\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\xi}^{3}}{n}\rightarrow 0,$ (2.9)
as $n\rightarrow\infty$, the following hold:
[…]
(ii) $\mathbb{W}_{1}$ provides a consistent test.
[…]
3.   SAR with error spatial dependence
Having demonstrated our methods with a baseline model, we now consider (2.1) but relax the $i.i.d$ assumption on $\epsilon_{i}$ to
$\epsilon_{i}\equiv\epsilon_{in}=\sum_{j=1}^{n}b_{ij}v_{j},$ (3.1)
where the $v_{j}$ are $i.i.d$ with mean 0 and variance 1, and the $b_{ij}\equiv b_{ijn}$ can depend on $n$. This is the Kelejian and Prucha (2007) approach to error spatial dependence, and indeed we will employ the spatial heteroskedasticity and autocorrelation consistent (SHAC) method pioneered in that paper. Our imposition that $z_{i}$ be uncorrelated with $\epsilon_{i}$ now, and subsequently, requires that $z_{i}$ is uncorrelated with all the $v_{j}$, $j=1,\ldots,n$.
3.1   Test statistic
Let $\mathscr{D}_{2}=\frac{1}{n}R\mathcal{U}_{1}R^{\prime}$ with $\mathcal{U}_{1}=T_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}\Xi_{1}(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}$ and $\widehat{\mathscr{D}}_{2}=\frac{1}{n}R\widehat{\mathcal{U}}_{1}R^{\prime}$ with $\widehat{\mathcal{U}}_{1}=T_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}\hat{\Xi}_{1}(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}$, where $\hat{\Xi}_{1}$ is the Kelejian and Prucha (2007) SHAC estimate of $\Xi_{1}=\frac{1}{n}K_{1}^{\prime}\Sigma K_{1}$, with $\Sigma=E(\epsilon\epsilon^{\prime})$ being the error variance matrix. We still use the 2SLS estimator $\hat{\xi}$ and the test statistic is now
$\mathbb{W}_{2}=\frac{n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{2}^{-1}\hat{\alpha}-d_{\alpha}}{\sqrt{2d_{\alpha}}}.$ (3.2)
We introduce the following assumptions:
Assumption 6.
$\Sigma$ has 1.
Assumption 7.
$\sup_{i\geq 1}\sum_{j=1}^{n}\left|b_{ij}\right|+\sup_{j\geq 1}\sum_{i=1}^{n}\left|b_{ij}\right|<\infty$.
Assumption 8.
$\mathscr{K}(\cdot):\mathbb{R}\rightarrow[-1,1]$ is a kernel function that satisfies $\mathscr{K}(0)=1$, $\mathscr{K}(x)=\mathscr{K}(-x)$, $\mathscr{K}(x)=0$ for $|x|>1$ and $|\mathscr{K}(x)-1|\leq C|x|^{\varrho},|x|\leq 1$, for some $\varrho\geq 1$.
Assumption 7 restricts the spatial dependence in the errors to a manageable degree, see e.g. Kelejian and Prucha (2007), Delgado and Robinson (2015) and Gupta and Qu (2024) for similar assumptions. Assumption 8 is a standard assumption on kernels in the HAC setting, see e.g. Kelejian and Prucha (2007).
Now we introduce distance measures $d_{ij,m}=d_{ji,m},m=1,\ldots,M$. As in Kelejian and Prucha (2007), we allow for measurement errors and so, in the following, let $d^{*}_{ij,m}=d^{*}_{ji,m}\geq 0$ be the actual distance measures used in practice. Corresponding to each measure, assume that the researcher can select a distance $d_{m}>0$ satisfying $d_{m}\uparrow\infty$ as $n\rightarrow\infty$. For each unit $i=1,\ldots,n$, let $\ell_{i}=\sum_{j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)$ and set $\ell=\max_{i}\ell_{i}$. Observe that $\ell_{i}$ is the number of units $j$ for which $d^{*}_{ij,m}\leq d_{m}$ for at least one $m=1,\ldots,M$. Also let $\sigma_{ij}$ be a typical element of $\Sigma$.
Assumption 9.
(a) $\mathrm{E}(\ell^{2})=o\left(n^{2\eta}\right)$ where $\eta<\frac{1}{2}(q-2)/(q-1)$ with $q>4$ in Assumption 2. (b) $\sum_{j=1}^{n}\left|\sigma_{ij}\right|d^{\chi}_{ij,1}<C$ for some $\chi\geq 1$. (c) $d^{*}_{ij,m}=d_{ij,m}+\nu_{ij,m}\geq 0$, with $\left|\nu_{ij,m}\right|<C$ and $\nu_{ij,m}$ independent of $v_{i}$ for all $m=1,\ldots,M$.
Then, the $(r,s)$-th element of $\hat{\Xi}_{1}$ is
$\hat{\Xi}_{1,rs}=n^{-1}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{ri}k_{sj}\hat{u}_{j}\hat{u}_{j}\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right),r,s=1,\ldots,J_{1}.$ (3.3)
3.2   Asymptotic properties
In all subsequent lemmas and theorems, $q$ is defined in Assumption 2 and $\eta$ is defined in Assumption 9. In Lemma 3.1 and Theorems 3.1-3.2, which follow, $d_{\xi}$ is defined as $d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$. Furthermore, for any matrix $A$, let $\|A\|=\left\{\overline{\alpha}\left(A^{\prime}A\right)\right\}^{\frac{1}{2}}$ i.e. the spectral norm of $A$.
Lemma 3.1.
Let Assumptions 1, 3, 4, 6-9 hold, and
$\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+n^{\eta-1}d_{\xi}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\xi}\rightarrow 0,\text{ as }n\rightarrow\infty.$ (3.4)
Then,
$\big{\|}\hat{\Xi}_{1}-\Xi_{1}\big{\|}=o_{p}(1).$
The condition $n^{\eta-1}d_{\xi}^{2}\rightarrow 0$ implies $d_{\xi}/n\rightarrow 0$ by Assumption 9(a). Now, define $\mathcal{M}_{2}=\frac{1}{n}B^{\prime}K_{1}\mathcal{V}_{2}K_{1}^{\prime}B$, with $\mathcal{V}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$. Then we have the following theorem.
Theorem 3.1.
Let (3.4) hold. Then, under $H_{0}$, Assumptions 1, 3, 4, 6-9, as $n\rightarrow\infty$,
$\mathbb{W}_{2}-\frac{v^{\prime}\mathcal{M}_{2}v-d_{\alpha}}{\sqrt{2d_{\alpha}}}=o_{p}(1).$
To establish the asymptotic distribution of $\mathbb{W}_{2}$, the following assumption is imposed to check a Lyapunov condition, as also e.g. by Delgado and Robinson (2015) and Gupta and Qu (2024).
Assumption 10.
$\mathrm{E}\left|v_{i}\right|^{s}<C$, for some $s\geq 8$.
Define $H_{\ell 2}\equiv H_{\ell 2,n}:\alpha=\alpha^{*}\equiv\nu_{2n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{2n}^{\prime}\Gamma_{2n}\nu_{2n})^{\frac{1}{2}}$, with $\nu_{2n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{2n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in the proof of the next theorem.
Theorem 3.2.
Under Assumptions 1, 3-10, with
$\frac{1}{d_{\lambda}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\xi}^{3}}{n}+n^{\eta-1}d_{\xi}^{2}+\ n^{\eta(q-1)/q-(q-2)/2q}d_{\xi}\rightarrow 0,$ (3.5)
as $n\rightarrow\infty$, the following hold:
[…]
(ii) $\mathbb{W}_{2}$ provides a consistent test.
[…]
Observe that the condition $n^{\eta-1}d_{\xi}^{2}\rightarrow 0$ implies $\frac{d_{\xi}^{3}}{n}\rightarrow 0$ if $\eta\geq 1/3$, so depending on the value of $q$, the rate condition (3.5) can be simplified.
4.   Misspecification robustness
We now consider a degree of misspecification robustness in (2.1), which allows $\epsilon_{i}$ to have the structure in Section 3. Focusing on the case $d_{\lambda}=1$ for notational simplicity, and writing $W_{1}=W$, (2.1) imposed the parametric form $\lambda Wy$ and took $W$ as known in the “spatial lag” term. We now allow the spatial lag in outcomes to take the nonparametric form $Gy$, where $G$ has elements $g\left(d_{rs}\right)$, $r,s=1,\ldots,n$, for some vector of exogenous (independent of $v_{j}$, $j=1,\ldots,n$) economic distance measures $d_{rs}$. This idea was introduced by Pinkse et al. (2002), and has been much used since (see e.g. Sun, 2016 and Gupta and Qu, 2024).
4.1   Test statistic
We approximate the elements $g\left(d_{ij}\right)$ of $G$ with a series of basis functions. Following Pinkse et al. (2002), we suppose
$g(d_{ij})=\sum_{l=1}^{\infty}\tau_{l}e_{l}(d_{ij}),$ (4.1)
where $\tau_{l}$ are unknown coefficients, and the $e_{l}$ form a basis of the function space to which $g$ belongs. Let $c_{i}^{\prime}$ be the $i$-th row of the $n\times d_{\tau}$ matrix $\mathfrak{C}$ with typical $(i,l)$-th element $\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $\tau=\left(\tau_{1},\ldots,\tau_{d_{\tau}}\right)^{\prime}$, $d_{\tau}\rightarrow\infty$ as $n\rightarrow\infty$. Then we generalize (2.1) to
$y_{in}=g_{in}^{\prime}y+x_{in}^{\prime}\beta+p_{in}^{\prime}\delta\left(z_{in}\right)+\epsilon_{in},i=1,\ldots,n,$ (4.2)
and (2.3) to
$y_{i}=c_{i}^{\prime}\tau+x_{i}^{\prime}\beta+\psi_{i}^{\prime}\alpha+u_{i},i=1,\ldots,n,$ (4.3)
with $u_{i}=r_{iG}+r_{i\delta}+\epsilon_{i}$, where $r_{iG}=\sum_{l=d_{\tau}+1}^{\infty}\tau_{l}\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $r_{i\delta}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$. In the matrix notation, we now have
$y=\mathfrak{C}\tau+X\beta+\Psi\alpha+u=F\theta+u,$ (4.4)
where $F=[\mathfrak{C},X,\Psi]$ and $\theta=\left(\tau^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$. Our test is implemented using the 2SLS estimator
$\hat{\theta}=\left(F^{\prime}{P_{K_{2}}}F\right)^{-1}F^{\prime}{P_{K_{2}}}y,$ (4.5)
where $P_{K_{2}}=K_{2}\left(K_{2}^{\prime}K_{2}\right)^{-1}K_{2}^{\prime}$, $K_{2}$ being an $n\times J_{2}$ instrument matrix with $J_{2}\geq d_{\theta}=d_{\tau}+d_{\beta}+d_{\alpha}$, but with $J_{2}$ and $d_{\theta}$ having the same asymptotic order.
Let $\mathscr{D}_{3}=\frac{1}{n}R\mathcal{U}_{2}R^{\prime}$ with $\mathcal{U}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\Xi_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$ and $T_{2}=\frac{1}{n}F^{\prime}{P_{K_{2}}}F$, and let $\widehat{\mathscr{D}}_{3}=\frac{1}{n}R\widehat{\mathcal{U}}_{2}R^{\prime}$ with $\widehat{\mathcal{U}}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\hat{\Xi}_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$, where $R=\left[0_{d_{\alpha}\times\left(d_{\tau}+d_{\beta}\right)},I_{d_{\alpha}}\right]$ and $\hat{\Xi}_{2}$ is the Kelejian and Prucha (2007) estimate of $\Xi_{2}=\frac{1}{n}K_{2}^{\prime}\Sigma K_{2}$, with $\Sigma=E(\epsilon\epsilon^{\prime})$. Then, the test statistic is
$\mathbb{W}_{3}=\frac{n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{3}^{-1}\hat{\alpha}-d_{\alpha}}{\sqrt{2d_{\alpha}}}.$ (4.6)
Denote the elements of $F$ as $f_{ri}$. We have following assumptions.
Assumption 11.
$\mathrm{E}\left(f^{2}_{ri}\right)<C$ and $\mathrm{E}\left(k_{ri}^{2}\right)<C$.
Assumption 12.
$\sup_{i\geq 1}\mathrm{E}\left(r_{iG}^{2}\right)=o(n^{-1})$ and $\sup_{i\geq 1}\mathrm{E}\left(r_{i\delta}^{2}\right)=o(n^{-1})$.
Assumption 13.
$n^{-1}K_{2}^{\prime}K_{2}$ and $n^{-1}F^{\prime}K_{2}K_{2}^{\prime}F$ have 1.
Lemma 4.1.
Let Assumptions 6-9, 11-13 hold, and
$\frac{1}{d_{\tau}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+n^{\eta-1}d_{\theta}^{2}+\ n^{\eta(q-1)/q-(q-2)/2q}d_{\theta}\rightarrow 0,\text{as }n\rightarrow\infty.$ (4.7)
Then,
$\bigg{\|}\hat{\Xi}_{2}-\Xi_{2}\bigg{\|}=o_{p}(1).$
4.2   Asymptotic properties
Define $\mathcal{M}_{3}=\frac{1}{n}B^{\prime}K_{2}\mathcal{V}_{3}K_{2}^{\prime}B$, with $\mathcal{V}_{3}=(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}R^{\prime}\mathscr{D}_{3}^{-1}RT_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}$. Then, we have the following theorem.
Theorem 4.1.
Let (4.7) hold. Then, under $H_{0}$, Assumptions 6-9, 11-13, as $n\rightarrow\infty$,
$\mathbb{W}_{3}-\frac{v^{\prime}\mathcal{M}_{3}v-d_{\alpha}}{\sqrt{2d_{\alpha}}}=o_{p}(1).$
Define $H_{\ell 3}\equiv H_{\ell 3,n}:\alpha=\alpha^{*}\equiv\nu_{3n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{3n}^{\prime}\Gamma_{3n}\nu_{3n})^{\frac{1}{2}}$, with $\nu_{3n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{3n}$ a $d_{\alpha}\times d_{\alpha}$ matrix. The following assumption generalizes Assumption 5 to the setting of this section.
Assumption 14.
$\left(I_{n}-G\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$, almost surely, uniformly in $d_{rs},r,s=1,\ldots,n$.
Theorem 4.2.
Under Assumptions 6-14, with
$\frac{1}{d_{\tau}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\theta}^{3}}{n}+n^{\eta-1}d_{\theta}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\theta}\rightarrow 0,$ (4.8)
as $n\rightarrow\infty$, the following hold:
[…]
(2) $\mathbb{W}_{3}$ provides a consistent test.
[…]
5.   Varying spatial coefficients
In this section we revert to the case of known $W_{j},j=1,\ldots,d_{\lambda}$, but now allow the spatial lag parameters $\lambda_{j}$ to vary. The model we consider is closely related to Malikov and Sun (2017) and Sun and Malikov (2018), with the difference that we do not study panel data settings. Instead, we allow $d_{\beta}\rightarrow\infty$ and general spatial dependence in the disturbances. Our model is also related to the one studied by Sun (2024), wherein the model includes a nonparametric nuisance regression function but does not feature dependent disturbances. Thus, our model complements the extant literature.
5.1   Test statistic
Specifically, we have
$y_{i}=\sum_{j=1}^{d_{\lambda}}\lambda_{j}\left(z_{i}\right)w_{i,j}^{\prime}y+x_{i}^{\prime}\beta+p_{i}^{\prime}\delta\left(z_{i}\right)+\epsilon_{i},i=1,\ldots,n,$ (5.1)
with $\lambda_{j}(\cdot)$ unknown scalar-valued functions and $\epsilon_{i}=\sum_{j=1}^{n}b_{ij}v_{j}$, i.e. error with spatial dependence. Introduce the approximation $\lambda_{j}\left(z_{i}\right)=\sum_{k=1}^{l_{j}}\mu_{kj}\phi_{kj}\left(z_{i}\right)+r_{ij\lambda},j=1,\ldots,d_{\lambda}$. Set $d_{\mu}=\sum_{j=1}^{d_{\lambda}}l_{j}$, $\phi_{j}=\left(\phi_{1j},\ldots,\phi_{l_{j}j}\right)^{\prime}$, and let $H(z)$ be the $n\times d_{\mu}$ matrix with $i$-th row $\left(y^{\prime}w_{i,1}\phi_{1}\left(z_{i}\right),\ldots,y^{\prime}w_{i,d_{\lambda}}\phi_{d_{\lambda}}\left(z_{i}\right)\right)$.
Then,
$y=G(z)\gamma+u,$ (5.2)
where $G(z)=\left[H(z),X,\Psi\right]$, $\gamma=\left(\mu^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$, $\mu=\left(\mu_{1}^{\prime},\ldots,\mu_{d_{\mu}}^{\prime}\right)$, and $u$ has $i$-th element $u_{i}=r_{i\lambda}+r_{i\delta}+\epsilon_{i}$ with $r_{i\lambda}=\sum_{j=1}^{d_{\lambda}}r_{ij\lambda}w_{i,j}^{\prime}y$. Defining $\Lambda_{j}(z)=diag\left(\lambda_{j}\left(z_{1}\right),\ldots,\lambda_{j}\left(z_{n}\right)\right)$ and $S(z)=I-\sum_{j=1}^{d_{\lambda}}\Lambda_{j}(z)W_{j}$, under the usual invertibility conditions,
$y=S(z)^{-1}\left(X\beta+t(z)+\epsilon\right),$ (5.3)
with $P$ having $i$-th row $p_{i}^{\prime}$ and $t(z)$ serving to stack $p_{i}^{\prime}\delta\left(z_{i}\right)$ into an $n\times 1$ vector with conformable subscript to $y_{i}$ and $\epsilon_{i}$.
The null hypothesis in this setting is
$H_{0}:\vartheta=0,$ (5.4)
where $\vartheta$ is either $\mu$ or $\alpha$ and our test will be implemented using the 2SLS estimator
$\hat{\gamma}=\left(G^{\prime}{P_{K_{3}}}G\right)^{-1}G^{\prime}{P_{K_{3}}}y,$ (5.5)
where $P_{K_{3}}=K_{3}\left(K_{3}^{\prime}K_{3}\right)^{-1}K_{3}^{\prime}$ with $K_{3}$ being an $n\times J_{3}$ instrument matrix with $J_{3}\geq d_{\gamma}=d_{\mu}+d_{\beta}+d_{\alpha}$, but with $J_{3}$ and $d_{\gamma}$ having the same asymptotic order. Note that we have omitted the $z$ argument for brevity.
Let $\mathscr{D}_{4}=\frac{1}{n}R\mathcal{U}_{3}R^{\prime}$ with $\mathcal{U}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\Xi_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$ and $T_{3}=\frac{1}{n}G^{\prime}{P_{K_{3}}}G$, and let $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$ with $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$, where $R=\left[I_{d_{\mu}},0_{d_{\mu}\times\left(d_{\beta}+d_{\alpha}\right)}\right]$ or $R=\left[0_{d_{\alpha}\times\left(d_{\mu}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $\hat{\Xi}_{3}$ is the Kelejian and Prucha (2007) estimate of $\Xi_{3}=\frac{1}{n}K_{3}^{\prime}\Sigma K_{3}$, with $\Sigma=E(\epsilon\epsilon^{\prime})$. Then, the test statistic is
$\mathbb{W}_{4}=\frac{n\hat{\vartheta}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\hat{\vartheta}-d_{\vartheta}}{\sqrt{2d_{\vartheta}}},$ (5.6)
where $d_{\vartheta}$ is either $d_{\mu}$ or $d_{\alpha}$.
Denote the elements of $G$ as $g_{ri}(z)$. We have the following assumptions:
Assumption 15.
$\sup_{z\in\mathcal{Z}}\mathrm{E}\left(g^{2}_{ri}(z)\right)<C$ and $\mathrm{E}\left(k_{ri}^{2}\right)<C$.
Assumption 16.
$\sup_{i\geq 1}\mathrm{E}\left(r_{i\lambda}^{2}\right)=o(n^{-1})$ and $\sup_{i\geq 1}\mathrm{E}\left(r_{i\delta}^{2}\right)=o(n^{-1})$.
Assumption 17.
$n^{-1}K_{3}^{\prime}K_{3}$ and $\sup_{z}n^{-1}G(z)^{\prime}K_{3}K_{3}^{\prime}G(z)$ have 1.
Lemma 5.1.
Let Assumptions 6-9, and 15 -17 hold, and
$\frac{1}{d_{\mu}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+n^{\eta-1}d_{\gamma}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\rightarrow 0,\text{ as }n\rightarrow\infty.$ (5.7)
Then, we have
$\bigg{\|}\hat{\Xi}_{3}-\Xi_{3}\bigg{\|}=o_{p}(1).$
5.2   Asymptotic properties
Define $\mathcal{M}_{4}=\frac{1}{n}B^{\prime}K_{3}\mathcal{V}_{4}K_{3}^{\prime}B$, with $\mathcal{V}_{4}=(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}R^{\prime}\mathscr{D}_{4}^{-1}RT_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}$. Then, we have the following theorem.
Theorem 5.1.
Let (5.7) hold. Then, under $H_{0}$, Assumptions 6-9, 15-17, as $n\rightarrow\infty$,
$\mathbb{W}_{4}-\frac{v^{\prime}\mathcal{M}_{4}v-d_{\vartheta}}{\sqrt{2d_{\vartheta}}}=o_{p}(1).$
The following assumption is the appropriate version of Assumptions 5 and 14 for the model considered in this section.
Assumption 18.
$\left(I_{n}-S(z)\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$, almost surely, uniformly in $z$.
Define $H_{\ell 4}\equiv H_{\ell 4,n}:\vartheta=\vartheta^{*}\equiv\nu_{4n}d_{\vartheta}^{\frac{1}{4}}/(n\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n})^{\frac{1}{2}}$, with $\nu_{4n}$ a $d_{\vartheta}\times 1$ non-zero vector and $\Gamma_{4n}$ a $d_{\vartheta}\times d_{\vartheta}$ matrix defined in the proof of the next theorem.
Theorem 5.2.
Under Assumptions 6-10 and 15-18, suppose that
$\frac{1}{d_{\mu}}+\frac{1}{d_{\beta}}+\frac{1}{d_{\alpha}}+\frac{d_{\gamma}^{3}}{n}+n^{\eta-1}d_{\gamma}^{2}+n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\rightarrow 0,$ (5.8)
as $n\rightarrow\infty$, the following hold:
[…]
(ii) $\mathbb{W}_{4}$ provides a consistent test.
[…]
6.   Empirical applications
6.1   Testing for the CRS in a production function
In this subsection, we are interested in studying whether the production function in China’s nonmetal mineral manufacturing industry has a constant returns to scale (CRS) technology. CRS refers to a scenario where a simultaneous proportional increase in all inputs results in an identical proportional increase in output. Returns to scale is a classical concept in economics to assess the efficiency of a production function, and it has accordingly received much attention in the empirical literature (for recent works, see Ackerberg et al., 2015; Basu and Bundick, 2017; Attanasio et al., 2020; Combes et al., 2021). Existing works, however, mostly do not account for cross-sectional dependence between firms. The role of cross-firms correlation has recently been highlighted empirically by Iyoha (2023), who shows that this can capture spillover effects and would lead to model misspecification if left unaccounted for.
We consider models that generalize Li et al. (2002)’s production function by incorporating spatial dependence in the analysis, thereby allowing for cross-sectional dependence. Their analysis centers on a Cobb-Douglas style production function that allows the elasticities of capital and labor to be nonparametrically varying with management expenses. Management expenses are costs that are indirectly associated with output production, such as research and development (R&D), equipment upgrades and employee training. In this framework, CRS corresponds to the elasticities of capital and labor summing to one.
Using data from the Third Industrial Census of China (conducted by the National Statistical Bureau in 1995) on the non-metal mineral manufacturing industry, Li et al. (2002) find that ignoring the varying nature of these coefficients leads to an underestimation of the returns to scale. Moreover, they suggest their estimate of the returns to scale of the technology is close to being constant for most values of managerial expense, but they do not have a test for the CRS hypothesis.
Our data comes from the Chinese Industrial Enterprise Database for the year of 2014 as this is the most recent information available in the database. We use data on all firms in the nonmetal mineral manufacturing industry with the survey code 31, resulting in a sample of 7,355 observations. As advocated by Li et al. (2002), the appeal for focusing on this industry is that production functions of firms are expected to be homogeneous since it has a very small proportion of foreign entity ownership. This is important because evidence suggests that production performance may vary across ownership types (Murakami et al., 1994). Table 1 introduces all variables after taking logarithms.
Our variables are output ($y_{i}$), capital ($p_{i1}$), labor ($p_{i2}$), and management expenses ($z_{i}$). As in Li et al. (2002), we include a parametric benchmark where the Cobb-Douglas production function has a quadratic form of management expenses without the varying coefficient entering additively, namely:
$y_{i}=\delta_{0}+\delta_{1}p_{i1}+\delta_{2}p_{i2}+\beta_{1}z_{i}+\beta_{2}z_{i}^{2}+\epsilon_{i}.$ (6.1)
The elasticities of capital and labor are $\delta_{1}$ and $\delta_{2}$, respectively. The semiparametric production function where the elasticities of of capital and labor vary with management expenses is:
$y_{i}=\delta_{0}(z_{i})+\delta_{1}(z_{i})p_{i1}+\delta_{2}(z_{i})p_{i2}+\epsilon_{i}.$
Based on the CRS hypothesis, $H_{0}^{true}:\delta_{1}(z_{i})+\delta_{2}(z_{i})=1$, we reparameterize the model as
$y_{i}^{*}=\delta_{0}(z_{i})+\delta_{2}(z_{i})(p_{i2}-p_{i1})+(\delta_{1}(z_{i})+\delta_{2}(z_{i})-1)p_{i1}+\epsilon_{i},$
where $y_{i}^{\ast}=y_{i}-p_{i1}$. Then, we estimate
$\begin{split}&(\mathbb{W}_{0})\ y_{i}^{*}=\beta_{0}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$ (6.2)
where $\psi_{ik}$ is a basis function, with argument $z_{i}$, for $k=1,...,h/2$. We approximate the unknown functions by polynomial and trigonometric basis functions of different orders. We define the series approximation based Wald statistic derived from this model with no spatial autoregressive component as $\mathbb{W}_{0}$.
Next, we incorporate the SAR structure and consider:
$y_{i}^{*}=\lambda(z_{i})\sum\limits_{j=1}^{n}w_{ij}y_{j}+\beta_{0}+\delta_{0}(z_{i})+\delta_{2}(z_{i})(p_{i2}-p_{i1})+(\delta_{1}(z_{i})+\delta_{2}(z_{i})-1)p_{i1}+\epsilon_{i}.$
We estimate a number of models and perform tests based on the various configurations and statistics discussed in previous sections:
$\begin{split}&(\mathbb{W}_{1})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{2})\ y_{i}^{*}=\beta_{0}+\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{3})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\\ &(\mathbb{W}_{4})\ y_{i}^{*}=\beta_{0}+\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=1}^{h}\alpha_{1k}\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{2k}(p_{i2}-p_{i1})\psi_{ik}+\sum\limits_{k=1}^{h}\alpha_{3k}p_{i1}\psi_{ik}+\epsilon_{i},\end{split}$ (6.3)
with $\epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j}$ for $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$. We use two different row-normalized spatial weight matrices, denoted $W_{n}^{p}$ and $W_{n}^{d}$:
$\begin{split}&w_{ij}=[W_{n}^{p}]_{ij}=\begin{cases}1\quad\text{iffirms}\ i\ \text{and}\ j\ \text{are in the same prefecture-level city},\\ 0\quad\text{otherwise},\end{cases}\\ &w_{ij}=[W_{n}^{d}]_{ij}=\begin{cases}\frac{1}{d_{ij,m}^{*}}\quad\text{if}\ d_{ij,m}^{*}<d_{m},\\ 0\quad\text{otherwise},\end{cases}\end{split}$
where $d_{ij,m}^{\ast}$ measures the geographical distance between firms $i$ and $j$, and $d_{m}$ represents the 10th percentile of all geographical distances.
For $\mathbb{W}_{i}$, $i=0,...,3$, $\alpha_{3}=(\alpha_{31},...\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha_{3}=0$. For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha_{3}=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$. Throughout our empirical analysis, we use $\psi_{ik}(z)=z_{i}^{k}$ as the polynomial function for $k=1,...,h$, and $\psi_{ik}(z)=[\sin(kz_{i}),\cos(kz_{i})]$ as the trigonometric function $k=1,...,h/2$ for $h=2,4$. To compute $\mathbb{W}_{2}$, $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$, we use the Epanechnikov kernel function to estimate $\mathscr{K}\left({d_{ij,m}^{\ast}}/{d_{m}}\right)$. For $\mathbb{W}_{3}$, we denote $\widehat{w}_{ij,l}=[d_{ij,m}^{\ast}]^{l}\mathbf{1}(d_{ij,m}^{\ast}<d_{m})$ if $i\neq j$ for $l=1,...,h$. Then, $c_{il}=\sum_{j=1}^{n}\widehat{w}_{ij,l}y_{j}$ for $l=1,...,h$. For $\mathbb{W}_{4}$, we use $\phi_{il}(z)=[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\sin(\frac{z_{i}}{2l})$ as the trigonometric function for $l=1,...,h$. Then, $h_{il}=y_{i}\sum_{j=1}^{n}w_{ij}\phi_{jl}(z)$. For the IVs, we use $K_{1}=[W\Psi,X,\Psi]$ with respect to $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$, and $K_{2}=[X,\Psi]$ and $K_{3}=[H\Psi,X,\Psi]$ with respect to $\mathbb{W}_{3}$ and $\mathbb{W}_{4}$.
The results from estimating the parametric model can be found in Table 2, and the results for other models with different basis functions are in Tables 3-6. In the latter group of tables, we define $\mathscr{T}_{n}$ as the asymptotic p-values corresponding to $\mathbb{W}_{i}$, $i=0,...,4$. It is well known that nonparametric tests, which rely on their asymptotic normal null distributions, may not perform well with finite samples. As an alternative, we provide bootstrap p-values that are denoted by $\mathscr{T}_{n}^{\ast}$ basing on 150 bootstrap replications. The details of the bootstrap procedure are given in Section 7.1.
Our results show that, in all models, the hypothesis of CRS cannot be rejected. These findings corroborate those in Li et al. (2002). A conclusion one can draw from this is that CRS is a salient feature of the production function in this industry, as it is present in the parametric model as well as in semiparametric models with varying coefficients and cross-firm dependence.
We end this application with some remarks that, in addition to inference on returns to scale, our analysis also reveals other similarities with prior works. Notably, from Table 7, we observe that the parametric method underestimates the returns to scale compared to the semiparametric method. [^1] The returns to scale in a semiparametric model can be computed using $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{1}(z_{i})$ and $\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{2}(z_{i})$. The returns to scale with respect to different input factors also differ, which highlights the crucial role of management expenses. To illustrate this, take the semiparametric estimates with spatial weights $W_{n}^{d}$ and produce Figure 1 and Figure 2. The former shows that the output elasticity of capital is increasing in management expenses, and the latter shows the output elasticity of labor is decreasing in management expenses.
Intuitively, by the nature of management expenses in production, the elasticities may be expected to increase with both inputs. However, the decreasing elasticities of labor is a known empirical fact in this industry due to the practice of “concealed unemployment”  where the government did not allow state-owned firms to lay off extra employees as a strategy to avoid social unrest. We also note that the returns to scale is relatively flat in the semiparametric model and is within the 95% bounds for most $z_{i}$, as illustrated in Figure 3. Further discussions on these findings can be found in Li et al. (2002). The results we have described here are not unique to the estimates from $W_{n}^{d}$. Other choices of spatial weights give the same conclusion; we simply selected one to illustrate our findings.
6.2   Testing the impact of the distance to the employment center on housing price
In this subsection, we use the Boston house price dataset, which has been used in several spatial studies, to illustrate another application of our proposed test. The dataset [^2] The related shp file can be downloaded at https://github.com/simonbrewer/geog6000/blob/main/datafiles/boston.tr.zip. Data descriptions are in Table 8. consists of the median value of owner-occupied homes in 506 census tracts in the Boston Standard Metropolitan Statistical Area in 1970 (Harrison Jr and Rubinfeld, 1978). Specifically, we explore how some factors such as the per capita crime rate by town (denoted by CRIM), average number of rooms per dwelling (RM), full-value property-tax rate per \$10,000 dollar (TAX), percentage of lower socioeconomic status population (LSTAT) and index of accessibility to radial highways (RAD) affect the median value of owner-occupied homes in $1000$s (MEDV), and whether the effects of these factors vary over location or not. DIS, which represents the weighted distances between a property and the five Boston employment centers that are associated with higher housing prices due to closer proximity to the city center, will drive the varying coefficient.
The motivation for considering spatial autocorrelation in house prices is straightforward, as housing prices in a specific location or region are commonly influenced by the prices of nearby houses due to factors such as environmental and geographic considerations. [^3] There are also some studies that use the SAR semiparametric model with the Boston house price data, for example, Malikov and Sun (2017) and Li et al. (2019). However, the former considers the covariate $z_{i}$ as NO2, which is more about the impact of environmental pollution on housing prices. The latter concentrates on the geographical weighted regression (GWR) method. Sun et al. (2014) proposed a semiparametric coefficient varying spatial model and derived the distribution theory of their estimator. While they do not have inference results on functions of the random coefficients, they have a model selection procedure to determine the parametric and nonparametric components in their model. We thus take the model they selected as our starting point.
Let $y_{i},x_{i1},x_{i2},p_{i1},p_{i2},p_{i3},z_{i}$ denote the log of MEDV, RAD, LSTAT, CRIM, RM, TAX and DIS respectively. Guided by model selection results, Sun et al. (2014) argue that DIS doesn’t affect LSTAT and RAD. Intuitively, for LSTAT, the geographical distribution of employment opportunities does not necessarily correlate with a community’s socio-economic status; a community might be far from employment centers but could have good public transportation links or a high number of remote work opportunities, mitigating the impact on the community’s lower socioeconomic status. For RAD, the accessibility to highways might depend more on urban planning and infrastructure investment rather than the proximity to employment centers.
The baseline model, based on Sun et al. (2014), is:
$y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i}.$
We focus on a more general version of their model, by allowing the spatial lag coefficient to be varying with DIS as well:
$y_{i}=\lambda(z_{i})\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i}.$
Therefore, we estimate:
$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{2}\\ &y_{i}=\sum\limits_{j=1}^{n}g(d_{ij})y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{3};\\ &y_{i}=\lambda(z_{i})\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\delta_{m}(z_{i})p_{im}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{4},\end{split}$
and the estimated models are
$\begin{split}&y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{and using}\ \mathbb{W}_{1};\\ &y_{i}=\lambda\sum\limits_{j=1}^{n}w_{ij}y_{j}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{2};\\ &y_{i}=\sum\limits_{l=1}^{h}\tau_{l}c_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{3};\\ &y_{i}=\sum\limits_{l=1}^{h}\mu_{l}h_{il}+\sum\limits_{k=0}^{2}\beta_{k}x_{ik}+\sum\limits_{m=1}^{3}\sum\limits_{k=1}^{h}\alpha_{mk}p_{im}\psi_{ik}+\epsilon_{i},\ \text{with}\ \epsilon_{i}=\sum\limits_{j=1}^{n}b_{ij}v_{j},\ \text{and using}\ \mathbb{W}_{4}.\end{split}$ (6.4)
We consider two row-normalized spatial weight matrices. The first is the first-order queen matrix, $W_{n}^{q}$, discussed in Malikov and Sun (2017). The second, denoted by $W_{n}^{t}$, indicates whether two locations are in the same tract. For $\mathbb{W}_{i}$, $i=0,...,3$, $\alpha=(\alpha_{11},...\alpha_{1h},...,\alpha_{3h})^{\prime}$, and the null is $H_{0}:\alpha=0$. For $\mathbb{W}_{4}$, the null is $H_{0}:\alpha=0$ or $H_{0}:\mu=0$, where $\mu=(\mu_{1},...,\mu_{h})^{\prime}$.
From Tables 9-12, we find that $log(DIS)$ indeed affects $p_{i1}$, $p_{i2}$ and $p_{i3}$, and has a significantly nonlinear impact on neighboring house prices. All test statistics show a rejection of the null $H_{0}:\alpha=0$. The bootstrap p-value also confirms this conclusion. In addition, testing with $\mathbb{W}_{4}$ also leads to a rejection of the null $H_{0}:\mu=0$, indicating a significant nonlinear relationship between $log(DIS)$ and neighboring house prices. Furthermore, we estimate the varying $\lambda(z_{i})$ and evaluate its empirical mean $\hat{\lambda}=\frac{1}{n}\sum_{i=1}^{n}\lambda(z_{i})$. From Table 13, we thus find that the parametric method underestimates the spatial coefficient. Our results, which are based on more general models, thus complement those in Sun et al. (2014).
7.   Monte Carlo Simulation
We present a basic set of Monte Carlo results in the main body of the paper. Because our theory covers such a wide range of models, which include several important special cases, we have performed a wide range of simulations for them too. However, the sheer number of tables we have makes it impractical to include them all in the main body of the paper, and we provide references to an online supplementary appendix.
7.1   Basic setting
Taking $n=100,300,500,$ we choose two specifications to generate $y$ with $d_{\delta}=1$:
$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i}\delta\left(z_{i}\right)+\epsilon_{i},i=1,\ldots,n,$ (7.1)
and
$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i}\delta\left(z_{i}\right)+v_{i},\ v_{i}=\gamma_{1}\sum\limits_{j=1}^{n}w_{2ij}v_{j}+\epsilon_{i},i=1,\ldots,n,$ (7.2)
[…]
Kelejian and Prucha (2007) consider the kernel function with a distance measure $d_{ij,m}^{*}$ and a threshold $d_{m}$. In estimation, we use the Epanechnikov kernel function, we choose the Euclidean distance $d_{ij}$ mentioned in $W_{1}$ to represent $d_{ij,m}^{*}$, and $d_{m}$ is considered as the quantile value of 10% of all distances. If $d_{ij,m}^{*}>d_{m}$, $\mathscr{K}\left(\frac{d_{ij,m}^{*}}{d_{m}}\right)=0$. We also consider $d_{\lambda}=2$ with $W_{1}$, $W_{2}$, and $\lambda=(0.5,0.3)^{\prime}$. The IV matrix for $\mathbb{W}_{1}$ and $\mathbb{W}_{2}$ is $K_{1}=[W_{1}X,W_{2}X,X,\Psi]$. We calculate the empirical rejection frequencies using the test statistics when nominal levels are given by 1%, 5% and 10%. We consider 1000 Monte Carlo replications for all experiments and $B=150$ bootstrap replications. We define $\mathscr{S}_{n}$ as the test statistic corresponding to $\mathbb{W}_{i}$ for $i=1,\ldots,4$ and $\mathscr{S}^{\ast}_{n}$ as the bootstrap version of $\mathscr{S}_{n}$.
The bootstrap version of our test, e.g., focusing on the setting described in Section 3, can be carried out by taking these steps:
  • Compute $\widehat{\epsilon}=y-L\hat{\xi}$ under null or alternative and obtain $\widetilde{\epsilon}=(I_{n}-\frac{1}{n}l_{n}l_{n}^{\prime})\widehat{\epsilon}$, where $l_{n}$ is the $n\times 1$ column vector of ones.
  • Sample randomly with replacement $n$ times from elements of $\widetilde{\epsilon}$ to obtain $\tilde{\epsilon}^{*}$.
  • Generate the bootstrap sample $\hat{y}=\left(I_{n}-\sum\limits_{j=1}^{d_{\lambda}}\hat{\lambda}_{j}W_{j}\right)^{+}\left([X,\Psi](\hat{\beta},\hat{\alpha})^{\prime}+\tilde{\epsilon}^{*}\right)$, where $(\cdot)^{+}$denotes the Moore-Penrose generalized inverse.
  • For the bootstrap, compute $\hat{\xi}^{\ast}$ and obtain the bootstrap residual $\hat{\epsilon}^{\ast}=\hat{y}-\hat{L}\hat{\xi}^{\ast}$ to construct $\mathscr{S}_{n}^{\ast}$, where $\hat{L}=[W_{1}\hat{y},\ldots,W_{d_{\lambda}}\hat{y},X,\Psi]$.
  • Repeat $B$ times to obtain the sequence of bootstrap test statistics $\left\{\mathscr{S}_{nj}^{\ast}\right\}_{j=1}^{B}$. Calculate bootstrap p-values $\mathscr{T}_{n}^{\ast}=\frac{1}{B}\sum\limits_{j=1}^{B}\mathbf{1}\left(\mathscr{S}_{n}<\mathscr{S}_{nj}^{\ast}\right)$ and reject the null hypothesis if $\mathscr{T}_{n}^{\ast}$ is smaller than the prescribed nominal level of significance.
As before, we denote by $\mathscr{T}_{n}$ the p-value based on the asymptotic test. From Tables 14-15, we find that the bootstrap test is closer to the nominal level than the asymptotic test although the sizes of both types of tests improve generally as the sample size increases. Specifically, we observe that the pairs $(n,h)=(500,8)$ deliver acceptable sizes under three disturbances. Moreover, we observe that the asymptotic test is powerful against local alternatives as $n$ increases from Tables S1-S2 in the supplement.
We also consider the case of $d_{\lambda}=2$ in the supplement. The size performances [^5] Power performances are close to or equal to 1 in most cases so we omit the power tables but they are available on request. are reported in Tables S3-S4 in the supplement and the finite sample properties are similar to the case of $d_{\lambda}=1$.
7.2   Nonparametric spatial matrix
For the nonparametric spatial matrix, we generate $w_{1ij}=\Phi\left(-b_{ij}\right)\mathbf{1}\left(c_{ij}<0.1\right)$ if $i\neq j$, and $w_{1ii}=0$, where $\Phi(\cdot)$ is the standard normal cdf, $b_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[-3,3]$, and $c_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U[0,1]$. From this construction, we ensure that $W_{1}$ is sparse with no more than $10\%$ elements being nonzero. Then, $W=W_{1}/1.2\bar{\alpha}(W_{1})$, ensuring the existence of $(I-W)^{-1}$. For $\mathbb{W}_{3}$, we approximate elements in $W$ by $\widehat{w}_{ij}=\sum_{l=0}^{r}\tau_{l}[d_{ij,m}^{*}]^{l}\mathbf{1}(d_{ij,m}^{*}<d_{m})$ if $i\neq j$ with $r=2,4$, where $d_{ij,m}^{*}$ and $d_{m}$ are discussed in Section 7.1, and the IV matrix for $\mathbb{W}_{3}$ is $K_{2}=[X,\Psi]$. We define this DGP setting as (D3).
Table 16 reports the simulation results at the 1%, 5% and 10% asymptotic level with $r=2,4$ and $h=2,4,8$. We find that there is a clear pattern of rejection rates approaching the theoretical level as sample size increases, and sizes are acceptable for $n=500$ and $h=8$. When $r$ increases, the results are similar. From Table S5, the power performances are close to or equal to 1 as $n$ increases. We also consider $W=W_{1}+W_{2}$ which corresponds to $d_{\lambda}=2$. Specifically, we set $w_{ij2}=d_{m}\exp(-d_{ij,m}^{*}/d_{m})$ if $d_{ij,m}^{*}\leq d_{m}$ and $w_{ij2}=0$ if $d_{ij,m}^{*}>d_{m}$. From Table S6 in the supplement, a similar pattern for size is observed.
7.3   Varying spatial coefficient
We now consider varying $\lambda$ such that $\lambda=\lambda(z_{i})$. When $d_{\lambda}=1$, we set $\lambda(z_{i})=\sin(0.75\pi\times z_{i})$, when $d_{\lambda}=2$, we set $\lambda_{1}(z_{i})=\sin(0.38\pi\times z_{i})$ and $\lambda_{2}(z_{i})=\cos(0.37\pi\times z_{i})$. Others are same as those in Section 7.1. For estimation, we use $\phi_{il}(z)=[\frac{2}{\pi}\tanh(z_{i})]^{l}$ as the polynomial function and $\phi_{il}(z)=\sin(\frac{z_{i}}{2l})$ as the trigonometric function for $l=1,...,h$. Then, $h_{1il}=y_{i}\sum_{j=1}^{n}w_{1ij}\phi_{jl}(z)$ and $h_{2il}=y_{i}\sum_{j=1}^{n}w_{2ij}\phi_{jl}(z)$. The IV matrix is $K_{3}=[H_{1}X,X,\Psi]$ for $d_{\lambda}=1$ and $K_{3}=[H_{1}X,H_{2}X,X,\Psi]$ for $d_{\lambda}=2$. We define this DGP setting as (D4).
Tables 17 and 18 report the empirical rejection frequencies using the test statistics when nominal levels are given by 1%, 5% and 10%. Sizes are acceptable as $n$ and $h$ increase, especially, the trigonometric approximation performs slightly better than polynomial approximations for Chi-square distribution when testing $\lambda=0$. Bootstrap test is closer to the nominal level than the asymptotic test although the sizes of both types of tests improve generally as the sample size increases. When $d_{\lambda}=2$, the results are similar, as shown in Tables S7-S8 in the supplement.
From Tables S9 and S10 in the supplement, power improves as $h$ increases for all designs, and the test for $\lambda$ and $\delta$ have similar results. As for $d_{\lambda}=2$, we find that our proposed test statistics still perform well as both $n$ and $h$ increase.
7.4   Additional simulation results in Online Supplement
In a range of experiments reported in the online supplement, we study how large $n$ with $h$ increasing to see how size changes. So we set $n=1000$ with $h=12,16$, $d_{\lambda}=1,2$ and $d_{\delta}=1$. From Table S11 of the supplement, we find that as $h$ increases, the performance of size improves.
We also consider tests of equality of varying coefficients and CRS hypothesis, where $d_{\delta}=2$ such that
$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\delta_{1}(z_{i})+p_{i2}\delta_{2}(z_{i})+\epsilon_{i},$ (7.3)
and
$y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\delta_{1}(z_{i})+p_{i2}\delta_{2}(z_{i})+v_{i},\ v_{i}=\gamma_{1}\sum\limits_{j=1}^{n}w_{2ij}v_{j}+\epsilon_{i},$ (7.4)
for $i=1,\ldots,n$. Our null of interest when testing equality is $H_{0}^{true}:\delta_{1}(z)=\delta_{2}(z),z\in\mathcal{Z},$ while for CRS it is $H_{0}^{true}:\delta_{1}(z)+\delta_{2}(z)=1,z\in\mathcal{Z}.$ In the first case, we reparameterize with $\theta(z)=\delta_{1}(z)-\delta_{2}(z)$, and the model is $y_{i}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\theta(z_{i})+(p_{i1}+p_{i2})\delta_{2}(z_{i})+\epsilon_{i}.$ In the second case, we reparameterize with $\theta(z)=\delta_{1}(z)+\delta_{2}(z)-1$ and the model is $y_{i}-p_{i1}=\lambda_{1}\sum\limits_{j=1}^{n}w_{1ij}y_{j}+x_{i}^{\prime}\beta+p_{i1}\theta(z_{i})+(p_{i2}-p_{i1})\delta_{2}(z_{i})+\epsilon_{i}.$
The results are included in the supplement. Taking CRS as an example, from Tables S12-S13, bootstrap tests are closer to the nominal levels than the asymptotic tests although the sizes of both types of tests improve generally as the sample size increases. For the nonparametric spatial matrix, from Table S14, there is a clear pattern of rejection rates approaching the theoretical level as sample size increases. The results of testing equality are similar, as shown in Tables S15-S17.
Appendix
Appendix A General Central Limit Theorem
Introduce an $n\times n$ matrix $\mathscr{A}$, which is symmetric, idempotent and has rank $J\rightarrow\infty$.
Assumption A.
Assume that
$\mathscr{A}=B^{\prime}\mathscr{S}B,$ (A.1)
where $B$ is the $n\times n$ matrix with entries $b_{ij}$ satisfying Assumption 7.
(b) The entries $s_{ij}$ of the symmetric matrix $\mathscr{S}$ satisfy $s_{ij}=O_{p}(J/n)$ and $\sum_{i=1}^{n}s_{ij}^{2}=O_{p}(J/n)$, uniformly in $i,j=1,\ldots,n$.
(c) $\mathscr{S}$ satisfies 1.
Theorem A.1.
Let $v$ be an $n\times 1$ $i.i.d.(0,1)$ vector with elements satisfying Assumption 10. Under Assumption A and $J^{-1}+J^{3}/n\rightarrow 0$ as $n\rightarrow\infty$,
[…]
Proof.
The proof is identical to that of Theorem 4.3 in Gupta and Qu (2024) and is omitted. ∎
Appendix B Proofs of Theorems
Proof of Theorem 2.1.
Denote $r_{\delta}=(r_{1\delta},\cdots,r_{n\delta})^{\prime}$. Note that $\hat{\xi}=\xi+(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}u=(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$ and $\hat{\alpha}=\alpha+R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$. Under $H_{0}:\alpha=0$, we have $\hat{\alpha}=R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)$. Then we are going to investigate
$\begin{split}&n(r_{\delta}+\epsilon)^{\prime}P_{K_{1}}L(L^{\prime}P_{K_{1}}L)^{-1}R^{\prime}\mathscr{D}_{1}^{-1}R(L^{\prime}P_{K_{1}}L)^{-1}L^{\prime}P_{K_{1}}(r_{\delta}+\epsilon)\\ =&\epsilon^{\prime}\mathcal{M}_{1}\epsilon+(A_{1}+A_{2})\end{split}$ (B.1)
where $\mathcal{M}_{1}=\frac{1}{n}K_{1}\mathcal{V}_{1}K_{1}^{\prime}$ with $\mathcal{V}_{1}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{1}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $T_{1}=\frac{1}{n}L^{\prime}{P_{K_{1}}}L$ and $\mathscr{D}_{1}=\frac{1}{n}RT_{1}^{-1}R^{\prime}$, and $A_{1}=r_{\delta}^{\prime}\mathcal{M}_{1}r_{\delta}$ and $A_{2}=2r_{\delta}^{\prime}\mathcal{M}_{1}\epsilon$. Now, $|A_{1}|\leq\|r_{\delta}\|^{2}\|\mathcal{M}_{1}\|=o_{p}(1)$, since $E\|r_{\delta}\|^{2}=o(1)$, and $\|\mathcal{M}_{1}\|=O_{p}(1)$ by Assumptions 1 and 4. A similar argument holds for $A_{2}$ because it has (conditionally) zero mean and variance $4r_{\delta}^{\prime}\mathcal{M}_{1}E(\epsilon\epsilon^{\prime})\mathcal{M}_{1}r_{\delta}=4r_{\delta}^{\prime}\mathcal{M}_{1}r_{\delta}$, by idempotence of $\mathcal{M}_{1}$. Thus $\frac{A_{1}}{\sqrt{d_{\alpha}}}$ and $\frac{A_{2}}{\sqrt{d_{\alpha}}}$ are both $o_{p}(1)$ and establishing the theorem. ∎
Proof of Theorem 2.2.
This is a special case of Theorem 3.2 and therefore omitted. ∎
Proof of Lemma 3.1.
This is a simpler version of the proof of Lemma 5.1 below and is therefore omitted. ∎
Proof of Theorem 3.1.
$\hat{\alpha}=\alpha+R\hat{\xi}=\alpha+R(L^{\prime}{P_{K_{1}}}L)^{-1}L^{\prime}{P_{K_{1}}}(r_{\delta}+Bv)$ and under $H_{0}:\alpha=0$,
$\begin{split}n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{2}^{-1}\hat{\alpha}=(v^{\prime}\widehat{\mathcal{M}}_{2}v+A_{11}+A_{12}+2A_{21}+2A_{22}),\end{split}$ (B.2)
where $A_{11}=r_{\delta}^{\prime}(\widehat{\mathscr{M}}_{2}-\mathscr{M}_{2})r_{\delta}$, $A_{12}=r_{\delta}^{\prime}\mathscr{M}_{2}r_{\delta}$, $A_{21}=r_{\delta}^{\prime}(\widehat{\mathcal{M}}_{2}-\mathcal{M}_{2})v$, $A_{22}=r_{\delta}^{\prime}\mathcal{M}_{2}v$, ^$\widehat{\mathcal{M}}_{2}=\frac{1}{n}B^{\prime}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}B$ with $\widehat{\mathcal{V}}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\widehat{\mathscr{D}}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$, $\widehat{\mathscr{D}}_{2}=\frac{1}{n}R\widehat{\mathcal{U}}_{1}R^{\prime}$, $\widehat{\mathcal{U}}_{1}=T_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}\hat{\Xi}_{1}(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}$, and ^$\widehat{\mathscr{M}}_{2}=\frac{1}{n}K_{1}\widehat{\mathcal{V}}_{2}K_{1}^{\prime}$. The remainder of the proof is similar to the proof of Theorem 5.1. ∎
Proof of Theorem 3.2.
The proof follows by setting $J=J_{1}$, $\mathscr{A}=\mathcal{M}_{2}$ and $\mathscr{S}=\frac{1}{n}K_{1}\mathcal{V}_{2}K_{1}^{\prime}$ in Theorem A.1. Checking Assumption A follows exactly as in the proof of Theorem 5.2 below and is omitted. ∎
Proof of Lemma 4.1.
This proof is a simpler version of the proof of Lemma 5.1 below and is therefore omitted. ∎
Proof of Theorem 4.1.
Define $r_{G}=(r_{1G},\cdots,r_{nG})^{\prime}$ and $r_{G\delta}=r_{G}+r_{\delta}$. Then, we have $\hat{\alpha}=\alpha+R\hat{\xi}=\alpha+R(F^{\prime}{P_{K_{2}}}F)^{-1}F^{\prime}{P_{K_{2}}}(r_{G\delta}+Bv)$ and $\big{\|}r_{G\delta}\big{\|}^{2}=o_{p}(1)$ by Assumption 12. Under $H_{0}:\alpha=0$,
$\begin{split}n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{3}^{-1}\hat{\alpha}=(v^{\prime}\widehat{\mathcal{M}}_{3}v+A_{11}+A_{12}+2A_{21}+2A_{22}),\end{split}$ (B.3)
where $A_{11}=r_{G\delta}^{\prime}(\widehat{\mathscr{M}}_{3}-\mathscr{M}_{3})r_{G\delta},A_{12}=r_{G\delta}^{\prime}\mathscr{M}_{3}r_{G\delta}$, $A_{21}=r_{G\delta}^{\prime}(\widehat{\mathcal{M}}_{3}-\mathcal{M}_{3})v$, $A_{22}=r_{G\delta}^{\prime}\mathcal{M}_{3}v$, ^$\widehat{\mathcal{M}}_{3}=\frac{1}{n}B^{\prime}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}B$ with $\widehat{\mathcal{V}}_{3}=(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}R^{\prime}\widehat{\mathscr{D}}_{3}^{-1}RT_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}$, $\widehat{\mathscr{D}}_{3}=\frac{1}{n}R\widehat{\mathcal{U}}_{2}R^{\prime}$, $\widehat{\mathcal{U}}_{2}=T_{2}^{-1}F^{\prime}K_{2}(K_{2}^{\prime}K_{2})^{-1}\allowbreak\hat{\Xi}_{2}(K_{2}^{\prime}K_{2})^{-1}K_{2}^{\prime}FT_{2}^{-1}$, and ^$\widehat{\mathscr{M}}_{3}=\frac{1}{n}K_{2}\widehat{\mathcal{V}}_{3}K_{2}^{\prime}$. The remainder of the proof is similar to the proof of Theorem 5.1. ∎
Proof of Theorem 4.2.
The proof is similar to that of Theorem 5.2 and omitted. ∎
Proof of Lemma 5.1.
Define $r_{\lambda}=(r_{1\lambda},...,r_{n\lambda})^{\prime}$. First, note that
$\hat{\gamma}-\gamma=(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}u=(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}(r_{\lambda}+r_{\delta}+\epsilon),$ (B.4)
and
$\begin{split}\mathrm{E}\left\|n^{-1}K_{3}^{\prime}(r_{\lambda}+r_{\delta})\right\|^{2}&\leq n^{-2}\left\|K_{3}\right\|^{2}\mathrm{E}\left\|r_{\lambda}+r_{\delta}\right\|^{2}\\ &\leq Cn^{-1}\overline{\alpha}(n^{-1}K_{3}^{\prime}K_{3})\left(\mathrm{E}\left\|r_{\lambda}\right\|^{2}+\mathrm{E}\left\|r_{\delta}\right\|^{2}\right)=o(1),\end{split}$ (B.5)
because $\mathrm{E}(r_{i\lambda}^{2})=o(n^{-1})$ and $\mathrm{E}(r_{i\delta}^{2})=o(n^{-1})$, by Assumptions 16 and 17.
Next,
$\mathrm{E}\left\|n^{-1}K_{3}^{\prime}\epsilon\right\|^{2}$$\leq$ $n^{-2}\sum_{i=1}^{n}\left\|k_{i}\right\|^{2}\mathrm{E}\epsilon_{i}^{2}+n^{-2}\sum_{i\neq j}\left\|k_{i}\right\|\left\|k_{j}\right\|\left|\mathrm{E}\epsilon_{i}\epsilon_{j}\right|$ (B.6)
$=$ $n^{-2}\sum_{i=1}^{n}\left\|k_{i}\right\|^{2}\sum_{r}b_{ir}^{2}+2n^{-2}\sum_{j}\left\|k_{j}\right\|\sum_{i<j}\left\|k_{i}\right\|\sup_{r}\left|b_{jr}\right|\left(\sum_{r}\left|b_{ir}\right|\right)$
$\leq$ $Cn^{-2}\sum_{i=1}^{n}\left\|k_{i}\right\|^{2}+Cn^{-2}\sup_{j}\left\|k_{j}\right\|\left(\sup_{r}\sum_{j}\left|b_{jr}\right|\right)\sum_{i<j}\left\|k_{i}\right\|$
$=$ $O(d_{\gamma}/n),$
by Assumptions 7 and 15.
Combining Assumption 17, (B.4), (B.5) and (B.6), we obtain
$\left\|\hat{\gamma}-\gamma\right\|=O_{p}(\sqrt{d_{\gamma}/n}).$ (B.7)
Now, observe that
$\hat{\Xi}_{r,s}-\Xi_{r,s}=a_{1,rs}+a_{2,rs}+a_{3,rs},$ (B.8)
where
$a_{1,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}k_{ir}k_{sj}(\hat{u}_{i}\hat{u}_{j}-\epsilon_{i}\epsilon_{j})\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right),$ (B.9)
$a_{2,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}k_{ir}k_{sj}(\epsilon_{i}\epsilon_{j}-\sigma_{ij})\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right),$ (B.10)
$a_{3,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}k_{ir}k_{sj}\sigma_{ij}\left(\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right)-1\right),$ (B.11)
where recall that $\sigma_{ij}$ is a typical element of $\Sigma$.
We will establish $a_{l,rs}=o_{p}(1),l=1,2,3$. To begin, write $r_{i}=r_{i\lambda}+r_{i\delta}$, $g_{i}^{\prime}$ for the $i$-th row of $G$ and note that
$\hat{u}_{i}\hat{u}_{j}-\epsilon_{i}\epsilon_{j}=g_{i}^{\prime}(\hat{\gamma}-\gamma)g_{j}^{\prime}(\hat{\gamma}-\gamma)+g_{i}^{\prime}(\hat{\gamma}-\gamma)r_{j}+r_{i}r_{j}+g_{j}^{\prime}(\hat{\gamma}-\gamma)\epsilon_{i}+g_{i}^{\prime}(\hat{\gamma}-\gamma)\epsilon_{j}+\epsilon_{i}r_{j}+\epsilon_{j}r_{i}.$ (B.12)
Now, by Assumptions 15 and 16, $\left|r_{i}\right|=o_{p}(1/\sqrt{n})$, $\left\|g_{i}\right\|=O_{p}(\sqrt{d_{\gamma}})$ and, as in the proof of Theorem 2 of Kelejian and Prucha (2007), $\sum_{j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\leq\ell$, all uniformly in $i$.
Then, we can write $\left|a_{1,rs}\right|\leq\sum_{l=1}^{7}a_{1l,rs}$, where, also using (B.7),
$a_{11,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{i}\right\|\left\|g_{j}\right\|\left\|\hat{\gamma}-\gamma\right\|^{2}\leq C\ell\cdot O_{p}\left(\frac{d_{\gamma}^{2}}{n}\right),$
$a_{12,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{i}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|r_{j}\right|\leq C\ell\cdot O_{p}\left(\frac{d_{\gamma}}{n}\right),$
$a_{13,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left|r_{i}\right|\left|r_{j}\right|\leq C\ell\cdot O_{p}\left(\frac{1}{n}\right),$
$a_{14,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{j}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|\epsilon_{i}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{d_{\gamma}}{n^{1/2-1/q}}\right),$
$a_{15,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left\|g_{i}\right\|\left\|\hat{\gamma}-\gamma\right\|\left|\epsilon_{j}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{d_{\gamma}}{n^{1/2-1/q}}\right),$
$a_{16,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left|\epsilon_{i}\right|\left|r_{j}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{1}{n^{1/2-1/q}}\right),$
$a_{17,rs}$ $=$ $n^{-1}\sum_{i,j=1}^{n}\left(1-\prod_{m=1}^{M}\mathbf{1}(d^{*}_{ij,m}>d_{m})\right)\left|\epsilon_{j}\right|\left|r_{i}\right|\leq C\ell^{1-1/q}\cdot O_{p}\left(\frac{1}{n^{1/2-1/q}}\right),$
where the Hölder inequality is used to bound $a_{1j,rs},j\geq 4$, as in the proof of Theorem 2 of Kelejian and Prucha (2007). Therefore, we need $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}(1)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}(1)$ as $n\rightarrow\infty$. By Assumption 9(a), we have $\frac{\ell d_{\gamma}^{2}}{n}=o_{p}\left(n^{\eta-1}d_{\gamma}^{2}\right)$ and $\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}=o_{p}\left(n^{\eta(q-1)/q-(q-2)/2q}d_{\gamma}\right)$. We then have $a_{1,rs}=o_{p}(1)$ under the sufficient rate in (5.7) as $n\rightarrow\infty$.
$a_{2,rs}=o_{p}(1)$ and $a_{3,rs}=o_{p}(1)$ are established exactly as the negligibility of $b_{rs,n}$ and $c_{rs,n}$ is established in the proof of Theorem 1 of Kelejian and Prucha (2007). This concludes the proof. ∎
Proof of Theorem 5.1.
Denote $r_{\lambda}=(r_{1\lambda},\cdots,r_{n\lambda})^{\prime}$, $r_{\lambda\delta}=r_{\lambda}+r_{\delta}$. Thus, $\hat{\vartheta}=\vartheta+R\hat{\xi}=\vartheta+R(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}{P_{K_{3}}}(r_{\lambda\delta}+Bv)$ and $\big{\|}r_{\lambda\delta}\big{\|}^{2}=o(1)$ by Assumption 16. Under $H_{0}:\vartheta=0$,
$\begin{split}n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\hat{\alpha}=(v^{\prime}\widehat{\mathcal{M}}_{4}v+A_{11}+A_{12}+2A_{21}+2A_{22})\end{split}$ (B.13)
where $A_{11}=r_{\lambda\delta}^{\prime}(\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4})r_{\lambda\delta},A_{12}=r_{\lambda\delta}^{\prime}\mathscr{M}_{4}r_{\lambda\delta}$, $A_{21}=r_{\lambda\delta}^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v$, $A_{22}=r_{\lambda\delta}^{\prime}\mathcal{M}_{4}v$, ^$\widehat{\mathcal{M}}_{4}=\frac{1}{n}B^{\prime}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}B$ with $\widehat{\mathcal{V}}_{4}=(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}R^{\prime}\widehat{\mathscr{D}}_{4}^{-1}RT_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}$, $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$, $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\allowbreak\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}$, and ^$\widehat{\mathscr{M}}_{4}=\frac{1}{n}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}$. Observe that $\mathcal{U}_{3}=T_{3}^{-1}\left(\frac{1}{n}L^{\prime}K_{3}\right)\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\Xi_{3}\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\left(\frac{1}{n}K_{3}^{\prime}L\right)T_{3}^{-1}$, we have $\big{\|}\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big{\|}\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ according to the proof of Lemma 5.1 and Assumptions 15 and 17. Thus, $\big{\|}\widehat{\mathcal{U}}_{3}^{-1}-\mathcal{U}_{3}^{-1}\big{\|}\leq\big{\|}\widehat{\mathcal{U}}_{3}^{-1}\big{\|}\big{\|}\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big{\|}\big{\|}\mathcal{U}_{3}^{-1}\big{\|}\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ by Assumptions 6 and 17, where $r_{n}=\frac{\ell d_{\gamma}^{2}}{n}+\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}$. Similarly, we have ^$\big{\|}\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4}\big{\|}=o_{p}(1)$, and ^$\big{\|}\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4}\big{\|}=o_{p}(1)$ by the structure of $\mathscr{M}_{4}$, $\mathcal{M}_{4}$ and Assumptions 7 and 15.
Thus, we have
$\frac{1}{\sqrt{d_{\vartheta}}}\big{|}v^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v\big{|}\leq Cr_{n}\cdot O_{p}\left(d_{\vartheta}^{-1/2}\right)=o_{p}(1)$ (B.14)
since $\mathrm{E}(v_{i}^{2})=1$.
Similar to (B.14) and the proof of Theorem 2.2, we have $\frac{1}{\sqrt{d_{\alpha}}}\left|A_{11}\right|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$, $\frac{1}{\sqrt{d_{\alpha}}}|A_{12}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ since $\mathscr{M}_{4}=O_{p}(1)$ by Assumptions 7 and 15, $\frac{1}{\sqrt{d_{\alpha}}}|A_{21}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ and $\frac{1}{\sqrt{d_{\alpha}}}|A_{22}|=o_{p}(1)$. Therefore, we obtain the sufficient rate from (5.7) as $n\rightarrow\infty$ for the required quantities to be negligible.
Proof of Theorem 5.2.
(i) Set $J=J_{3}$, $\mathscr{A}=\mathcal{M}_{4}$ and $\mathscr{S}=\frac{1}{n}K_{3}\mathcal{V}_{4}K_{3}^{\prime}$ in Theorem A.1. We now proceed as in the proof Theorem 4.3 of Gupta and Qu (2024) and note that the entries of $\mathscr{S}$ are,
$s_{ij,4}=\frac{1}{n}k_{i}^{\prime}\mathcal{V}_{4}k_{j},$ (B.15)
where $k_{i}^{\prime}$ is the $i$-th row of $K_{3}$, and so
$\left|s_{ij,4}\right|=O_{p}\left(n^{-1}\left\|k_{i}\right\|\left\|k_{j}\right\|\right)=O_{p}\left(d_{\gamma}n^{-1}\right),$ (B.16)
uniformly in $i,j$, by Assumptions 15 and 17, and arguments used in the proof of Theorem 5.1 above.
Similarly, we also observe that
$\sum_{j=1}^{n}s_{ij,4}^{2}=O_{p}\left(d_{\gamma}n^{-1}\right),$ (B.17)
uniformly in $i$. Thus Assumption A(b) is satisfied. Assumption A is satisfied to due to Assumption 17. The conclusion now follows by Theorem A.1.
(ii) This proof is similar to that of Theorem 4.4 in Gupta and Qu (2024).
(iii) We prove all local alternatives here. We define $H_{\ell i}\equiv H_{\ell i,n}$, with the corresponding statistic being $\mathbb{W}_{i}^{*}$, for $i=1,2,3,4$. For $\mathbb{W}_{1}^{*}$, we consider $\alpha=\alpha^{*}\equiv\nu_{1n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{\frac{1}{2}}$ with $\nu_{1n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{1n}$ a $d_{\alpha}\times d_{\alpha}$ matrix. Under $H_{\ell 1}$, we have
$\mathbb{W}_{1}^{*}=\mathbb{W}_{1}+\frac{1}{\sqrt{2d_{\alpha}}}d_{\alpha}^{\frac{1}{2}}\mathcal{A}_{1}(\frac{1}{n}\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{-1}+o_{p}(1),$
where $\mathcal{A}_{1}=\frac{1}{n}\nu_{1n}^{\prime}\mathscr{D}_{1}^{-1}\nu_{1n}$. We choose $\Gamma_{1n}=\mathscr{D}_{1}^{-1}$ and then finish the proof. For $\mathbb{W}_{4}^{*}$, we consider $\vartheta=\vartheta^{*}\equiv\nu_{4n}d_{\vartheta}^{\frac{1}{4}}/(n\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n})^{\frac{1}{2}}$. Under $H_{\ell 4}$, we have
$\mathbb{W}_{4}^{*}=\mathbb{W}_{4}+\frac{1}{\sqrt{2d_{\vartheta}}}d_{\vartheta}^{\frac{1}{2}}\mathcal{A}_{4}(\frac{1}{n}\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n})^{-1}+o_{p}(1),$
where $\mathcal{A}_{4}=\frac{1}{n}\nu_{4n}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\nu_{4n}$, and $\big{\|}\frac{1}{n}\nu_{4n}^{\prime}(\widehat{\mathscr{D}}_{4}^{-1}-\mathscr{D}_{4}^{-1})\nu_{4n}\big{\|}=o_{p}(1)$ by (5.7). We choose $\Gamma_{4n}=\widehat{\mathscr{D}}_{4}^{-1}$ and then finish the proof. The proofs of $\mathbb{W}_{2}^{*}$ and $\mathbb{W}_{3}^{*}$ are similar. ∎
Appendix C Empirical results
Table 1: Data descriptions in Section 6.1, logarithm form.
Variables Descriptions Mean Min Max Sd
$y$ Value added 11.539 3.611 19.216 1.847
$p_{1}$ Capital assets 10.280 1.609 19.160 1.980
$p_{2}$ Average number of employ 5.647 2.021 11.865 1.130
$z$ Management expense 7.707 2.303 16.139 1.853
Table 2: Results of baseline model according to (6.1) in Section 6.1.
                 $\delta_{0}$                  3.313***
                 $\delta_{1}$                  0.554***
                 $\delta_{2}$                  0.402***
                 $\beta_{1}$                  0.224
                 $\beta_{2}$                  0.018
                 $H_{0}:\delta_{1}+\delta_{2}-1=0$
                 t-test                  -0.238
                 p-value                  0.812
                 $n$                  7355
                 The standard errors of the estimators
                 are reported in parentheses.
                  $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 3: Estimation and testing results in Section 6.1 using polynomial series with $h=2$.
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
$\lambda/\tau_{1}/\mu_{1}$ 0.189* 0.130** 0.224*** 0.256*** 0.451*** 0.457*** 0.499***
(0.112) (0.062) (0.014) (0.081) (0.157) (0.042) (0.028)
$\tau_{2}/\mu_{2}$0.095 0.405*** 0.484***
(0.126) (0.047) (0.030)
$\beta_{0}$ 1.073** 0.167* 0.153** 0.167*** 0.153*** -0.485*** -0.093*** 0.077***
(0.514) (0.100) (0.076) (0.016) (0.020) (0.013) (0.004) (0.003)
$\alpha_{11}$ 1.254** -0.036*** -0.032*** -0.036 -0.032 -0.023 0.005 0.004
(0.551) (0.007) (0.004) (0.027) (0.035) (0.975) (0.033) (0.026)
$\alpha_{12}$ 0.375 -0.015 0.017 -0.015 0.017 -0.032 0.015*** 0.027***
(3.048) (0.035) (0.019) (0.010) (0.036) (0.050) (0.002) (0.002)
$\alpha_{21}$0.194 -0.004 -0.004 -0.004 -0.004 -0.002 0.000 0.003
(0.786) (0.020) (0.033) (0.005) (0.012) (0.111) (0.001) (0.002)
$\alpha_{22}$0.023 0.092*** 0.088*** 0.092*** 0.088*** 0.032*** -0.017*** 0.012***
(0.077) (0.013) (0.013) (0.004) (0.029) (0.008) (0.000) (0.000)
$\alpha_{31}$0.046 0.031 0.019 0.031 0.019* 0.011 0.001 0.003
(0.891) (0.017) (0.010) (0.022) (0.011) (0.136) (0.018) (0.014)
$\alpha_{32}$0.062 0.004 0.005 0.004 0.005 0.059 -0.033 0.026
(0.092) (0.023) (0.023) (0.001) (0.004) (0.060) (0.031) (0.051)
$\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha_{3}=0$ 1.593 3.311 3.280 1.419 4.020 0.226 1.716 2.589
$\mathscr{T}_{n}$ 0.451 0.191 0.194 0.492 0.134 0.893 0.424 0.274
$\mathscr{T}_{n}^{*}$ 0.767 0.213 0.167 0.413 0.167 0.867 0.433 0.213
$H_{0}:\mu=0$ 68.365 92.106
$\mathscr{T}_{n}$ $<$0.001 $<$0.001
$\mathscr{T}_{n}^{*}$ 0.007 0.007
$n$ 7355
The standard errors of estimators are reported in parentheses. $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 4: Estimation and testing results in Section 6.1 using trigonometric series with $h=2$.
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
$\lambda/\tau_{1}/\mu_{1}$ 0.065 0.162* 0.065** 0.162*** 0.582*** 0.607*** 0.297***
(0.052) (0.087) (0.033) (0.014) (0.098) (0.166) (0.088)
$\tau_{2}/\mu_{2}$0.276*** 0.147** 0.494***
(0.041) (0.074) (0.088)
$\beta_{0}$ 1.111*** -0.381*** -0.382** -0.381*** -0.382*** -0.187*** 0.747*** 1.076***
(0.020) (0.136) (0.162) (0.042) (0.021) (0.012) (0.245) (0.380)
$\alpha_{11}$ 0.949*** 0.167*** 0.160*** 0.167*** 0.160*** 0.125*** -0.253*** -0.024***
(0.150) (0.023) (0.025) (0.014) (0.012) (0.016) (0.025) (0.004)
$\alpha_{12}$ 0.861*** -0.049 -0.028** -0.049 -0.028 -0.124*** 0.146*** 0.034***
(0.151) (0.033) (0.012) (0.035) (0.032) (0.008) (0.040) (0.007)
$\alpha_{21}$ 0.001 0.0096 0.094 0.096*** 0.094*** 0.014*** -0.091*** 0.021***
(0.007) (0.112) (0.124) (0.003) (0.003) (0.002) (0.019) (0.005)
$\alpha_{22}$ 0.002 0.026 0.024 0.026*** 0.024*** -0.018*** 0.020*** 0.009***
(0.007) (0.069) (0.043) (0.004) (0.003) (0.002) (0.005) (0.003)
$\alpha_{31}$ 0.017 0.085 0.068 0.085 0.068 0.001 -0.020 0.019
(0.018) (0.246) (0.191) (0.105) (0.064) (0.001) (0.017) (0.012)
$\alpha_{32}$ 0.009 0.033 0.025 0.033 0.025 -0.003 0.047 0.021
(0.022) (0.090) (0.084) (0.002) (0.002) (0.003) (0.007) (0.006)
$\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha_{3}=0$ 0.777 1.178 0.877 1.086 0.943 1.575 1.707 2.169
$\mathscr{T}_{n}$ 0.678 0.555 0.645 0.581 0.624 0.455 0.426 0.338
$\mathscr{T}_{n}^{*}$ 0.660 0.533 0.680 0.553 0.600 0.467 0.267 0.333
$H_{0}:\mu=0$ 321.305 406.511
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.013 0.027
$n$ 7355
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 5: Estimation and testing results in Section 6.1 using polynomial series with $h=4$.
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
$\lambda/\tau_{1}/\mu_{1}$ 0.330*** 0.265** 0.330*** 0.065*** 0.444*** 0.236*** 0.313***
(0.095) (0.134) (0.019) (0.024) (0.135) (0.043) (0.0969)
$\tau_{2}/\mu_{2}$0.013 0.219*** 0.330***
(0.045) (0.052) (0.064)
$\tau_{3}/\mu_{3}$ 0.025** 0.203** 0.258**
(0.012) (0.082) (0.126)
$\tau_{4}/\mu_{4}$0.088 0.189* 0.096*
(0.068) (0.110) (0.051)
$\beta_{0}$ 1.866 0.040*** 0.005*** 0.040*** 0.005 -0.036 0.075 -0.278***
(4.659) (0.011) (0.001) (0.011) (0.023) (0.159) (0.221) (0.062)
$\alpha_{11}$ 1.471 0.002*** 0.000*** 0.002 0.000 0.000 0.003 -0.006
(4.108) (0.000) (0.000) (1.759) (2.686) (10.471) (26.990) (3.119)
$\alpha_{12}$1.067 0.005*** 0.002*** 0.005 0.002 0.010 0.009 -0.005
(5.407) (0.001) (0.000) (0.329) (0.588) (2.398) (5.235) (0.903)
$\alpha_{13}$ 1.186 0.013*** 0.007*** 0.013 0.007 0.035 0.014 0.010
(8.484) (0.001) (0.002) (0.028) (0.053) (0.232) (0.466) (0.100)
$\alpha_{14}$0.020 0.121*** -0.004 0.121*** -0.004*** -0.174*** 0.180*** -0.046***
(0.288) (0.021) (0.042) (0.001) (0.002) (0.008) (0.016) (0.004)
$\alpha_{21}$0.032 0.015*** 0.003*** 0.015 0.003 0.000 0.006 -0.004
(0.676) (0.004) (0.001) (0.180) (0.462) (0.861) (3.899) (0.134)
$\alpha_{22}$0.029 0.041*** 0.020*** 0.041 0.020 0.004 0.013 0.022
(2.100) (0.008) (0.006) (0.031) (0.100) (0.190) (0.642) (0.071)
$\alpha_{23}$ 0.003 -0.010 0.092*** -0.010*** 0.092*** -0.011 -0.018 0.019*
(0.213) (0.019) (0.026) (0.002) (0.009) (0.017) (0.046) (0.010)
$\alpha_{24}$ $0$0.005*** -0.003** -0.005*** -0.003*** 0.005*** -0.001 -0.001***
(0.007) (0.000) (0.001) (0.000) (0.000) (0.001) (0.001) (0.000)
$\alpha_{31}$ 0.005 -0.005 -0.002*** -0.005 -0.002 -0.001 -0.002 0.006
(0.012) (0.061) (0.001) (0.344) (0.556) (1.561) (3.200) (0.827)
$\alpha_{32}$0.009 -0.008 -0.009 -0.008 -0.009 -0.009 -0.002 0.004
(0.042) (0.042) (0.053) (0.068) (0.124) (0.428) (0.543) (0.297)
$\alpha_{33}$ 0.007 0.059 -0.041 0.059 -0.041 -0.032 0.029 0.031
(0.032) (0.096) (0.113) (0.046) (0.033) (0.041) (0.040) (0.031)
$\alpha_{34}$0.002 0.001 0.016*** 0.001 0.016 0.008 -0.004 -0.003
(0.026) (0.073) (0.157) (0.006) (0.062) (0.011) (0.083) (0.044)
$\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha_{3}=0$ 0.075 1.707 1.464 2.401 1.808 0.030 0.098 0.385
$\mathscr{T}_{n}$ 0.963 0.426 0.481 0.301 0.405 0.985 0.952 0.825
$\mathscr{T}_{n}^{*}$ 0.947 0.440 0.500 0.287 0.467 0.933 0.887 0.793
$H_{0}:\mu=0$ 101.231 122.398
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.007 0.007
$n$ 7355
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 6: Estimation and testing results in Section 6.1 using trigonometric series with $h=4$.
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
$\lambda/\tau_{1}/\mu_{1}$ 0.224*** 0.256*** 0.224*** 0.256*** 0.196** 0.123* 0.136***
(0.040) (0.083) (0.014) (0.081) (0.098) (0.067) (0.033)
$\tau_{2}/\mu_{2}$ 0.261* 0.298** 0.269**
(0.136) (0.143) (0.110)
$\tau_{3}/\mu_{3}$ 0.029* 0.293** 0.299**
(0.018) (0.145) (0.140)
$\tau_{4}/\mu_{4}$0.011 0.107* 0.138*
(0.124) 0.026** 0.035***
$\beta_{0}$ 1.108*** 0.060** 0.004** 0.060*** 0.004*** -0.067*** (0.008) (0.009)
(0.021) (0.031) (0.002) (0.009) (0.001) (0.025) (0.055) (0.077)
$\alpha_{11}$ 0.995*** 0.167*** 0.153*** 0.167*** 0.153*** -0.415*** 0.143*** 0.150***
(0.162) (0.010) (0.011) (0.016) (0.020) (0.014) (0.016) (0.014)
$\alpha_{12}$ 0.904*** -0.036*** -0.032*** -0.036 -0.032 0.735*** -0.032 -0.022
(0.164) (0.005) (0.006) (0.027) (0.035) (0.019) (0.023) (0.021)
$\alpha_{13}$ 0.084 -0.015** 0.017*** -0.015 0.017 -0.198*** -0.003 -0.026***
(0.158) (0.007) (0.005) (0.010) (0.036) (0.006) (0.008) (0.006)
$\alpha_{14}$0.251 -0.004*** -0.004 -0.004 -0.004 -0.160*** -0.002 -0.019***
(0.160) (0.001) (0.003) (0.005) (0.012) (0.007) (0.005) (0.004)
$\alpha_{21}$ 0.001 0.092*** 0.088*** 0.092*** 0.088*** -0.006*** 0.037*** 0.037***
(0.007) (0.015) (0.016) (0.004) (0.029) (0.001) (0.003) (0.003)
$\alpha_{22}$ 0.001 0.031*** 0.019*** 0.031*** 0.019*** -0.037*** 0.013*** 0.014***
(0.008) (0.012) (0.003) (0.002) (0.003) (0.003) (0.002) (0.002)
$\alpha_{23}$ 0.000 0.004 0.035 0.004*** 0.035*** 0.014*** 0.007*** -0.003***
(0.007) (0.021) (0.022) (0.001) (0.004) (0.001) (0.001) (0.001)
$\alpha_{24}$ 0.005 0.003 0.006 0.003*** 0.006*** 0.003*** 0.002*** 0.002***
(0.008) (0.004) (0.008) (0.000) (0.001) (0.001) (0.000) (0.000)
$\alpha_{31}$ 0.018 0.087 0.068 0.087 0.068 -0.027 0.041 0.041
(1.320) (1.065) (1.012) (0.127) (0.126) (0.863) (0.079) (0.166)
$\alpha_{32}$ 0.020 0.029 0.025 0.029 0.025 -0.081 0.013 0.006
(0.685) (2.392) (2.003) (0.268) (0.212) (0.062) (0.091) (0.099)
$\alpha_{33}$ 0.003 0.028 0.005 0.028 0.005 0.019 0.010 0.009
(0.331) (5.691) (4.001) (0.316) (0.305) (0.891) (1.063) (0.085)
$\alpha_{34}$ 0.005 -0.022 -0.020 -0.022 -0.020 0.019 -0.011 -0.013
(0.369) (0.355) (0.093) (0.287) (0.365) (1.605) (0.062) (0.076)
$\mathbb{W}_{0}$ $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha_{3}=0$ 0.290 1.848 1.497 0.038 1.443 0.034 0.375 1.022
$\mathscr{T}_{n}$ 0.865 0.397 0.473 0.981 0.486 0.983 0.829 0.600
$\mathscr{T}_{n}^{*}$ 0.873 0.393 0.453 0.953 0.467 0.927 0.833 0.580
$H_{0}:\mu=0$ 86.301 90.233
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.027 0.007
$n$ 7355
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 7: Comparison of parametric method and semiparametric method in Section 6.1.
No $W$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
parametric $\hat{\delta}_{1}$ 0.5536 0.5438 0.5330 0.5989 0.5996 0.5960
$\hat{\delta}_{2}$ 0.4022 0.3794 0.3725 0.3145 0.3226 0.3099
$\hat{\delta}_{1}$+$\hat{\delta}_{2}$ 0.9558 0.9232 0.9055 0.9134 0.9222 0.9060
polynomial, $h=2$ $\hat{\delta}_{1}$ 0.6228 0.6488 0.6179 0.6263 0.6051 0.6108
$\hat{\delta}_{2}$ 0.3664 0.3212 0.3691 0.2930 0.3502 0.3220
$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9892 0.9699 0.9869 0.9194 0.9553 0.9328
trigonometric,$h=2$ $\hat{\delta}_{1}$ 0.6233 0.6257 0.6149 0.6284 0.6298 0.6806
$\hat{\delta}_{2}$ 0.3517 0.3099 0.3381 0.3040 0.3377 0.2816
$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9750 0.9356 0.9530 0.9324 0.9675 0.9621
polynomial, $h=4$ $\hat{\delta}_{1}$ 0.6840 0.6390 0.5432 0.6077 0.6213 0.6342
$\hat{\delta}_{2}$ 0.3089 0.3395 0.4345 0.3252 0.3471 0.3025
$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9929 0.9785 0.9777 0.9330 0.9683 0.9368
trigonometric,$h=4$ $\hat{\delta}_{1}$ 0.6792 0.6896 0.6814 0.6152 0.6289 0.6569
$\hat{\delta}_{2}$ 0.3121 0.2702 0.3002 0.3043 0.3649 0.3199
$\hat{\delta}_{1}+\hat{\delta}_{2}$ 0.9913 0.9598 0.9816 0.9195 0.9938 0.9768
$n$ 7355
The result in the first column does not consider any spatial weight matrix.
Figure 1: Comparison of parametric method and semiparametric method in Section 6.1, capital.
Figure 2: Comparison of parametric method and semiparametric method in Section 6.1, labor.
Figure 3: Comparison of parametric method and semiparametric method in Section 6.1, CRS.
Table 8: Data descriptions in Section 6.2.
Variables Descriptions Mean Min Mean SD
$y$ log(MEDV) 3.035 1.609 3.912 0.408
$x_{1}$ log(LSTAT) 2.371 0.548 3.637 0.601
$x_{2}$ log(RAD) 1.868 0 3.178 0.875
$p_{1}$ log(CRIM) -0.780 -5.064 4.488 2.162
$p_{2}$ log(RM) 1.832 1.270 2.172 0.112
$p_{3}$ log(TAX) 5.931 5.231 6.567 0.396
$z$ log(DIS) 1.188 0.122 2.495 0.540
Table 9: Estimation and testing results in Section 6.2 using polynomial series with $h=2$.
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
$\lambda/\tau_{1}/\mu_{1}$ 0.238*** 0.219*** 0.238*** 0.219*** 0.214** 0.512*** 0.468***
(0.049) (0.044) (0.028) (0.022) (0.084) (0.073) (0.089)
$\tau_{2}/\mu_{2}$ 0.330** -0.356*** -0.311***
(0.129) (0.049) (0.057)
$\beta_{0}$ 4.081*** 3.987*** 4.081*** 3.987*** -0.003 0.132*** 0.022*
(0.267) (0.272) (0.147) (0.156) (0.006) (0.018) (0.012)
$\beta_{1}$0.432*** -0.432*** -0.432*** -0.432*** -0.143* -0.053*** -0.162
(0.025) (0.024) (0.027) (0.025) (0.075) (0.004) (0.175)
$\beta_{2}$ 0.166*** 0.154*** 0.166*** 0.154*** 0.086* 0.025*** 0.044**
(0.024) (0.023) (0.018) (0.017) (0.049) (0.001) (0.020)
$\alpha_{11}$0.044*** -0.041*** -0.044*** -0.041* -0.056* -0.033** -0.017***
(0.006) (0.006) (0.006) (0.024) (0.030) (0.001) (0.004)
$\alpha_{12}$ 0.028*** 0.025*** 0.028*** 0.025*** 0.043* 0.022*** 0.014***
(0.004) (0.004) (0.005) (0.005) (0.022) (0.001) (0.003)
$\alpha_{21}$ 0.042** 0.046** 0.042** 0.046*** 0.115** -0.001 -0.068***
(0.019) (0.018) (0.019) (0.006) (0.051) (0.002) (0.011)
$\alpha_{22}$0.012 -0.015 -0.012 -0.015 0.136*** 0.008*** 0.035***
(0.012) (0.011) (0.009) (0.022) (0.049) (0.001) (0.004)
$\alpha_{31}$$0**$027*** -$0**$027*** $0*$005*** -0.019***
(0.007) (0.007) (0.006) (0.002) (0.030) (0.002) (0.003)
$\alpha_{32}$ 0.076* 0.070* 0.076*** $0**$084*** 0.082*** 0.101***
(0.040) (0.040) (0.003) (0.004) (0.015) (0.021) (0.011)
Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha=0$ 40.041 33.913 290.165 33.913 183.122 91.459 123.439
$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.013 0.020 0.007 0.013 0.013 0.020 0.007
$H_{0}:\mu=0$ 32.336 35.141
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.007 0.013
$n$ 506
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 10: Estimation and testing results in Section 6.2 using trigonometric series with $h=2$.
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
$\lambda/\tau_{1}/\mu_{1}$ 0.254*** 0.225*** 0.254*** 0.225*** 0.259*** 0.177*** 0.121**
(0.049) (0.044) (0.028) (0.023) (0.069) (0.027) (0.057)
$\tau_{2}/\mu_{2}$ 0.397*** -0.256*** -0.204**
(0.149) (0.040) (0.083)
$\beta_{0}$ 3.527*** 3.547*** 3.527*** 3.547*** 0.001 0.567*** 0.675***
(0.236) (0.229) (0.175) (0.139) (0.009) (0.091) (0.028)
$\beta_{1}$0.410*** -0.415*** -0.410*** -0.415*** -0.235*** -0.396*** -0.456***
(0.024) (0.023) (0.023) (0.024) (0.064) (0.021) (0.032)
$\beta_{2}$ 0.156*** 0.146*** 0.156*** 0.146*** 0.170*** 0.154*** 0.170***
(0.024) (0.023) (0.017) (0.017) (0.031) (0.018) (0.020)
$\alpha_{11}$0.056* -0.035* -0.056*** -0.035*** -0.024* -0.006** -0.007*
(0.040) (0.021) (0.003) (0.003) (0.012) (0.003) (0.004)
$\alpha_{12}$0.023*** -0.022*** -0.023*** -0.022*** -0.163*** -$0**$024***
(0.005) (0.005) (0.006) (0.005) (0.027) (0.005) (0.005)
$\alpha_{21}$ 0.054*** 0.049*** 0.054*** 0.049*** 0.241** 0.037*** 0.048***
(0.010) (0.010) (0.008) (0.007) (0.097) (0.009) (0.011)
$\alpha_{22}$0.016 -0.007 -0.016*** -0.007*** 0.212*** 0.106** 0.132**
(0.018) (0.018) (0.003) (0.002) (0.026) (0.052) (0.065)
$\alpha_{31}$0.027*** -0.023*** -0.027*** -0.023*** -0.017 -0.025*** -0.030***
(0.004) (0.004) (0.002) (0.002) (0.036) (0.002) (0.002)
$\alpha_{32}$ 0.010* 0.009* 0.010*** 0.009** 0.622*** 0.053*** 0.065***
(0.006) (0.005) (0.004) (0.004) (0.015) (0.004) (0.005)
Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha=0$ 38.897 34.423 229.115 325.327 266.301 175.332 144.982
$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.020 0.007 0.007 0.020 0.007 0.013 0.007
$H_{0}:\mu=0$ 28.883 21.326
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.027 0.007
$n$ 506
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 11: Estimation and testing results in Section 6.2 using polynomial series with $h=4$.
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
$\lambda/\tau_{1}/\mu_{1}$ 0.239*** 0.269*** 0.239** 0.269*** 0.162** 0.169** 0.169***
(0.050) (0.043) (0.095) (0.027) (0.069) (0.080) (0.019)
$\tau_{2}/\mu_{2}$ 0.175*** -0.363** -0.363***
(0.042) (0.177) (0.080)
$\tau_{3}/\mu_{3}$ 0.241*** 0.345** 0.345***
(0.062) (0.147) (0.113)
$\tau_{4}/\mu_{4}$ 0.082** -0.122** -0.123**
(0.035) (0.060) (0.053)
$\beta_{0}$ 3.285*** 2.925*** 3.285 2.925*** -$0**$092* -0.440**
(0.464) (0.444) (7.140) (0.347) (0.003) (0.047) (0.211)
$\beta_{1}$0.399*** -0.386*** -0.399*** -0.386*** -0.516*** 0.021* -0.105***
(0.026) (0.025) (0.039) (0.025) (0.092) (0.012) (0.013)
$\beta_{2}$ 0.158*** 0.137*** 0.158*** 0.137*** 0.313*** 0.019* 0.052***
(0.025) (0.025) (0.023) (0.018) (0.057) (0.011) (0.013)
$\alpha_{11}$ 0.041 0.031 0.041 0.031 0.048* 0.034** 0.006**
(0.053) (0.050) (0.067) (0.039) (0.027) (0.015) (0.003)
$\alpha_{12}$0.224* -0.185* -0.224* -0.185** -0.120*** -0.072** -0.012*
(0.118) (0.112) (0.135) (0.086) (0.024) (0.034) (0.006)
$\alpha_{13}$ 0.228*** 0.193** 0.228*** 0.193*** -0.055* 0.050*** 0.080*
(0.087) (0.083) (0.087) (0.062) (0.028) (0.016) (0.046)
$\alpha_{14}$0.064*** -0.055*** -0.064*** -0.055*** 0.058*** -0.011** -0.018*
(0.021) (0.020) (0.018) (0.015) (0.018) (0.005) (0.010)
$\alpha_{21}$0.706*** -$0**$706*** -$0**$073 0.019 0.003**
(0.240) (0.226) (0.226) (0.114) (0.073) (0.011) (0.001)
$\alpha_{22}$ 1.574*** 1.596*** 1.574*** 1.596*** -0.014 -0.026* -0.006**
(0.484) (0.453) (0.460) (0.229) (0.045) (0.015) (0.002)
$\alpha_{23}$1.040*** -1.078*** -1.040*** -1.078*** -0.020 0.056*** 0.030***
(0.312) (0.291) (0.295) (0.147) (0.016) (0.014) (0.009)
$\alpha_{24}$ 0.217*** 0.230*** 0.217*** 0.230*** -0.057* $0**$045**
(0.065) (0.060) (0.060) (0.030) (0.030) (0.003) (0.018)
$\alpha_{31}$ 0.262*** 0.264*** 0.262*** 0.264*** 0.068*** -$0**$017***
(0.084) (0.079) (0.080) (0.041) (0.015) (0.003) (0.006)
$\alpha_{32}$0.572*** -0.577*** -0.572*** -0.577*** 0.032 0.019*** 0.035**
(0.166) (0.155) (0.164) (0.079) (0.027) (0.007) (0.014)
$\alpha_{33}$ 0.371*** 0.381*** 0.371*** 0.381*** -0.026 -0.089* -0.021***
(0.106) (0.099) (0.105) (0.050) (0.035) (0.046) (0.004)
$\alpha_{34}$0.076*** -$0**$076*** -0.080*** 0.029* 0.011* 0.038***
(0.022) (0.020) (0.021) (0.010) (0.015) (0.006) (0.007)
$\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha=0$ 33.876 20.486 392.020 928.892 601.967 394.776 352.386
$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.007 0.013 0.013 0.007 0.020 0.013 0.013
$H_{0}:\mu=0$ 56.124 66.981
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.027 0.007
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 12: Estimation and testing results in Section 6.2 using trigonometric series with $h=4$.
Matrices $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $\widehat{W}$ $W_{n}^{q}$ $W_{n}^{t}$
$\lambda/\tau_{1}/\mu_{1}$ 0.230*** 0.275*** 0.230*** 0.275*** 0.211*** 0.039 0.043***
(0.050) (0.043) (0.026) (0.031) (0.065) (0.284) (0.005)
$\tau_{2}/\mu_{2}$ 0.258*** -0.397** -0.437***
(0.039) (0.186) (0.048)
$\tau_{3}/\mu_{3}$ 0.255*** 0.229** 0.253***
(0.055) (0.107) (0.028)
$\tau_{4}/\mu_{4}$ 0.092*** 0.333* 0.367***
(0.031) (0.190) (0.041)
$\beta_{0}$ 3.207*** 2.762*** 3.207*** 2.762*** -0.005 0.001** 0.001***
(0.454) (0.432) (0.871) (0.290) (0.004) (0.001) (0.000)
$\beta_{1}$0.392*** -0.376*** -0.392*** -0.376*** -0.296*** -0.257** -0.200***
(0.026) (0.024) (0.031) (0.024) (0.061) (0.111) (0.025)
$\beta_{2}$ 0.146*** 0.124*** 0.146*** 0.124*** -0.046 0.103 0.069***
(0.025) (0.024) (0.019) (0.016) (0.036) (0.299) (0.010)
$\alpha_{11}$0.089*** -0.082*** -0.089*** -0.082*** 0.234*** -0.074** -0.060***
(0.027) (0.025) (0.019) (0.021) (0.078) (0.032) (0.017)
$\alpha_{12}$ 0.201*** 0.175*** 0.201*** 0.175*** 0.355*** 0.162** 0.068*
(0.068) (0.064) (0.046) (0.045) (0.093) (0.071) (0.038)
$\alpha_{13}$$0**$084*** -$0**$084*** -0.315*** -0.075** -0.020
(0.031) (0.029) (0.022) (0.019) (0.095) (0.033) (0.017)
$\alpha_{14}$$0**$081*** -$0**$081*** 0.236*** -0.073*** -0.054***
(0.026) (0.025) (0.019) (0.021) (0.066) (0.022) (0.017)
$\alpha_{21}$ 0.086*** 0.104*** 0.086*** 0.104*** -0.283*** 0.064*** 0.107***
(0.031) (0.029) (0.021) (0.016) (0.085) (0.019) (0.011)
$\alpha_{22}$0.611*** -0.688*** -0.611*** -0.688*** -0.005 -0.394 -0.506***
(0.179) (0.165) (0.170) (0.146) (0.062) (1.827) (0.138)
$\alpha_{23}$ 0.354*** 0.396*** 0.354*** 0.396*** 0.305*** 0.229*** 0.278***
(0.103) (0.095) (0.100) (0.086) (0.108) (0.071) (0.080)
$\alpha_{24}$ 0.025 0.050 0.025 $0**$042 0.036 0.091***
(0.036) (0.034) (0.019) (0.014) (0.033) (0.076) (0.011)
$\alpha_{31}$0.027** -0.027** -0.027*** -0.027*** 0.128*** -0.025*** -0.035***
(0.013) (0.012) (0.009) (0.007) (0.032) (0.006) (0.005)
$\alpha_{32}$ 0.216*** 0.238*** 0.216*** 0.238*** -0.138 0.150* 0.217***
(0.061) (0.056) (0.057) (0.048) (0.086) (0.088) (0.043)
$\alpha_{33}$0.123*** -0.136*** -0.123*** -0.136*** -0.047 -0.086** -0.121***
(0.035) (0.033) (0.033) (0.029) (0.046) (0.038) (0.025)
$\alpha_{34}$0.004 -0.012 -0.004 -0.012** -0.014 -0.009 -0.030***
(0.013) (0.012) (0.008) (0.005) (0.026) (0.018) (0.005)
Testing statistics $\mathbb{W}_{1}$ $\mathbb{W}_{1}$ $\mathbb{W}_{2}$ $\mathbb{W}_{2}$ $\mathbb{W}_{3}$ $\mathbb{W}_{4}$ $\mathbb{W}_{4}$
$H_{0}:\alpha=0$ 40.041 33.913 290.165 33.913 258.704 91.459 123.439
$\mathscr{T}_{n}$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.013 0.020 0.007 0.013 0.013 0.020 0.007
$H_{0}:\mu=0$ 43.155 46.397
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.027 0.007
$n$ 506
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 13: Comparison of parametric method and semiparametric method in Section 6.2.
polynomial,$h=2$ polynomial, $h=4$ trigonometric,$h=2$ trigonometric,$h=4$
$W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$ $W_{n}^{q}$ $W_{n}^{t}$
0.3136 0.3371 0.3252 0.3406 0.3074 0.3221 0.2996 0.3125
Appendix D Simulation results
Table 14: Empirical sizes for $\mathbb{W}_{1}$ with $d_{\lambda}=1$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level.
polynomial series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.044 0.072 0.102 0.038 0.072 0.094 0.048 0.076 0.092
$\mathscr{T}_{n}^{*}$ 0.029 0.078 0.088 0.029 0.069 0.118 0.019 0.078 0.127
$h=4$ $\mathscr{T}_{n}$ 0.044 0.090 0.118 0.032 0.080 0.110 0.036 0.082 0.110
$\mathscr{T}_{n}^{*}$ 0.010 0.029 0.088 0.078 0.127 0.137 0.019 0.108 0.147
$h=8$ $\mathscr{T}_{n}$ 0.058 0.132 0.184 0.032 0.106 0.146 0.038 0.064 0.094
$\mathscr{T}_{n}^{*}$ 0.049 0.088 0.137 0.028 0.127 0.137 0.019 0.098 0.108
7.1 $h=2$ $\mathscr{T}_{n}$ 0.046 0.100 0.132 0.026 0.076 0.110 0.048 0.082 0.110
$\mathscr{T}_{n}^{*}$ 0.029 0.137 0.167 0.020 0.088 0.098 0.020 0.059 0.118
$h=4$ $\mathscr{T}_{n}$ 0.048 0.100 0.142 0.024 0.066 0.100 0.026 0.066 0.100
$\mathscr{T}_{n}^{*}$ 0.078 0.147 0.186 0.019 0.078 0.118 0.019 0.078 0.137
$h=8$ $\mathscr{T}_{n}$ 0.060 0.120 0.178 0.042 0.074 0.120 0.018 0.054 0.104
$\mathscr{T}_{n}^{*}$ 0.038 0.176 0.196 0.039 0.078 0.147 0.020 0.078 0.137
7.1 $h=2$ $\mathscr{T}_{n}$ 0.062 0.108 0.140 0.028 0.058 0.092 0.026 0.056 0.098
$\mathscr{T}_{n}^{*}$ 0.029 0.069 0.098 0.008 0.020 0.039 0.019 0.088 0.118
$h=4$ $\mathscr{T}_{n}$ 0.072 0.116 0.170 0.020 0.066 0.116 0.036 0.078 0.108
$\mathscr{T}_{n}^{*}$ 0.049 0.078 0.127 0.029 0.069 0.088 0.015 0.108 0.147
$h=8$ $\mathscr{T}_{n}$ 0.094 0.168 0.216 0.026 0.062 0.112 0.034 0.078 0.136
$\mathscr{T}_{n}^{*}$ 0.058 0.127 0.186 0.029 0.108 0.167 0.014 0.088 0.137
trigonometric series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.040 0.064 0.098 0.044 0.070 0.104 0.046 0.070 0.092
$\mathscr{T}_{n}^{*}$ 0.029 0.059 0.078 0.029 0.039 0.088 0.019 0.048 0.118
$h=4$ $\mathscr{T}_{n}$ 0.052 0.092 0.122 0.030 0.076 0.140 0.032 0.074 0.116
$\mathscr{T}_{n}^{*}$ 0.010 0.029 0.069 0.039 0.108 0.118 0.019 0.058 0.137
$h=8$ $\mathscr{T}_{n}$ 0.068 0.138 0.180 0.038 0.110 0.134 0.034 0.066 0.094
$\mathscr{T}_{n}^{*}$ 0.029 0.078 0.137 0.029 0.118 0.157 0.019 0.038 0.108
7.1 $h=2$ $\mathscr{T}_{n}$ 0.044 0.084 0.112 0.038 0.084 0.118 0.048 0.084 0.114
$\mathscr{T}_{n}^{*}$ 0.078 0.147 0.176 0.029 0.059 0.108 0.010 0.039 0.088
$h=4$ $\mathscr{T}_{n}$ 0.050 0.108 0.152 0.022 0.066 0.104 0.032 0.070 0.114
$\mathscr{T}_{n}^{*}$ 0.069 0.118 0.186 0.049 0.098 0.127 0.019 0.058 0.127
$h=8$ $\mathscr{T}_{n}$ 0.064 0.108 0.162 0.032 0.086 0.124 0.022 0.052 0.080
$\mathscr{T}_{n}^{*}$ 0.069 0.127 0.176 0.020 0.108 0.157 0.019 0.058 0.157
7.1 $h=2$ $\mathscr{T}_{n}$ 0.060 0.090 0.120 0.034 0.070 0.112 0.022 0.062 0.098
$\mathscr{T}_{n}^{*}$ 0.039 0.069 0.098 0.020 0.039 0.078 0.018 0.068 0.137
$h=4$ $\mathscr{T}_{n}$ 0.064 0.094 0.164 0.022 0.076 0.114 0.032 0.070 0.106
$\mathscr{T}_{n}^{*}$ 0.078 0.098 0.147 0.039 0.088 0.098 0.017 0.068 0.118
$h=8$ $\mathscr{T}_{n}$ 0.098 0.158 0.206 0.024 0.060 0.102 0.032 0.078 0.120
$\mathscr{T}_{n}^{*}$ 0.069 0.127 0.176 0.039 0.127 0.167 0.016 0.068 0.127
Table 15: Empirical sizes for $\mathbb{W}_{2}$ with $d_{\lambda}=1$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level.
polynomial series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.088 0.200 0.248 0.094 0.156 0.212 0.090 0.190 0.266
$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.108 0.049 0.088 0.127 0.020 0.039 0.069
$h=4$ $\mathscr{T}_{n}$ 0.060 0.150 0.206 0.084 0.170 0.218 0.086 0.172 0.204
$\mathscr{T}_{n}^{*}$ 0.010 0.039 0.039 0.029 0.049 0.069 0.049 0.127 0.157
$h=8$ $\mathscr{T}_{n}$ 0.062 0.134 0.194 0.062 0.136 0.174 0.068 0.144 0.202
$\mathscr{T}_{n}^{*}$ 0.020 0.078 0.108 0.049 0.078 0.098 0.020 0.127 0.137
7.1 $h=2$ $\mathscr{T}_{n}$ 0.102 0.142 0.182 0.054 0.096 0.152 0.052 0.106 0.136
$\mathscr{T}_{n}^{*}$ 0.029 0.059 0.118 0.059 0.108 0.127 0.020 0.078 0.176
$h=4$ $\mathscr{T}_{n}$ 0.098 0.158 0.212 0.042 0.090 0.136 0.036 0.086 0.114
$\mathscr{T}_{n}^{*}$ 0.020 0.069 0.118 0.069 0.118 0.137 0.029 0.108 0.147
$h=8$ $\mathscr{T}_{n}$ 0.120 0.210 0.282 0.066 0.122 0.162 0.034 0.088 0.144
$\mathscr{T}_{n}^{*}$ 0.010 0.059 0.088 0.049 0.127 0.157 0.029 0.069 0.167
7.1 $h=2$ $\mathscr{T}_{n}$ 0.068 0.118 0.142 0.024 0.054 0.084 0.028 0.078 0.092
$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.078 0.000 0.020 0.088 0.069 0.069 0.108
$h=4$ $\mathscr{T}_{n}$ 0.054 0.106 0.164 0.024 0.062 0.098 0.048 0.076 0.108
$\mathscr{T}_{n}^{*}$ 0.039 0.069 0.069 0.020 0.059 0.088 0.039 0.098 0.127
$h=8$ $\mathscr{T}_{n}$ 0.062 0.138 0.174 0.028 0.058 0.092 0.032 0.084 0.132
$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.078 0.029 0.069 0.098 0.039 0.078 0.127
trigonometric series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.082 0.160 0.212 0.106 0.198 0.246 0.084 0.176 0.232
$\mathscr{T}_{n}^{*}$ 0.029 0.069 0.098 0.078 0.108 0.127 0.020 0.088 0.118
$h=4$ $\mathscr{T}_{n}$ 0.072 0.146 0.208 0.076 0.160 0.204 0.086 0.150 0.208
$\mathscr{T}_{n}^{*}$ 0.029 0.118 0.118 0.029 0.039 0.078 0.029 0.069 0.098
$h=8$ $\mathscr{T}_{n}$ 0.086 0.144 0.194 0.076 0.154 0.204 0.084 0.152 0.184
$\mathscr{T}_{n}^{*}$ 0.039 0.108 0.137 0.029 0.078 0.098 0.049 0.108 0.118
7.1 $h=2$ $\mathscr{T}_{n}$ 0.086 0.134 0.180 0.056 0.112 0.140 0.056 0.108 0.144
$\mathscr{T}_{n}^{*}$ 0.039 0.078 0.118 0.039 0.108 0.137 0.029 0.049 0.147
$h=4$ $\mathscr{T}_{n}$ 0.102 0.188 0.240 0.040 0.100 0.140 0.052 0.092 0.138
$\mathscr{T}_{n}^{*}$ 0.020 0.059 0.069 0.078 0.137 0.167 0.049 0.108 0.147
$h=8$ $\mathscr{T}_{n}$ 0.112 0.192 0.270 0.052 0.106 0.160 0.032 0.076 0.132
$\mathscr{T}_{n}^{*}$ 0.010 0.029 0.049 0.029 0.167 0.176 0.049 0.118 0.167
7.1 $h=2$ $\mathscr{T}_{n}$ 0.046 0.096 0.120 0.026 0.064 0.098 0.026 0.072 0.102
$\mathscr{T}_{n}^{*}$ 0.039 0.059 0.069 0.020 0.059 0.078 0.059 0.098 0.127
$h=4$ $\mathscr{T}_{n}$ 0.040 0.104 0.150 0.022 0.066 0.104 0.034 0.068 0.118
$\mathscr{T}_{n}^{*}$ 0.020 0.078 0.118 0.049 0.088 0.088 0.059 0.088 0.127
$h=8$ $\mathscr{T}_{n}$ 0.056 0.132 0.158 0.022 0.064 0.096 0.024 0.076 0.130
$\mathscr{T}_{n}^{*}$ 0.029 0.069 0.078 0.029 0.078 0.127 0.049 0.088 0.108
Table 16: Empirical sizes for $\mathbb{W}_{3}$ with $d_{\lambda}=1$, rejection probabilities $\mathscr{T}_{n}$ at 1%, 5% and 10% level.
polynomial series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
$r=2$ 7.1 $h=2$ 0.032 0.096 0.164 0.040 0.096 0.116 0.014 0.056 0.096
$h=4$ 0.028 0.056 0.084 0.024 0.052 0.084 0.036 0.060 0.080
$h=8$ 0.016 0.036 0.052 0.024 0.036 0.048 0.024 0.040 0.092
7.1 $h=2$ 0.040 0.088 0.148 0.044 0.096 0.120 0.020 0.060 0.112
$h=4$ 0.040 0.088 0.104 0.032 0.100 0.120 0.012 0.044 0.066
$h=8$ 0.016 0.048 0.064 0.008 0.016 0.028 0.028 0.056 0.092
7.1 $h=2$ 0.040 0.104 0.116 0.024 0.076 0.104 0.048 0.092 0.124
$h=4$ 0.028 0.052 0.088 0.032 0.048 0.068 0.024 0.072 0.104
$h=8$ 0.020 0.048 0.052 0.024 0.056 0.072 0.004 0.024 0.052
$r=4$ 7.1 $h=2$ 0.032 0.120 0.156 0.040 0.084 0.108 0.016 0.052 0.106
$h=4$ 0.036 0.064 0.092 0.032 0.076 0.088 0.020 0.040 0.076
$h=8$ 0.012 0.032 0.040 0.012 0.032 0.048 0.004 0.032 0.044
7.1 $h=2$ 0.032 0.100 0.136 0.032 0.076 0.092 0.048 0.100 0.144
$h=4$ 0.044 0.096 0.100 0.052 0.092 0.116 0.004 0.028 0.056
$h=8$ 0.008 0.044 0.056 0.008 0.036 0.052 0.024 0.052 0.080
7.1 $h=2$ 0.056 0.092 0.128 0.048 0.080 0.132 0.060 0.100 0.132
$h=4$ 0.028 0.060 0.084 0.008 0.044 0.056 0.020 0.068 0.120
$h=8$ 0.032 0.044 0.060 0.024 0.060 0.084 0.020 0.048 0.064
trigonometric series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
$r=2$ 7.1 $h=2$ 0.028 0.092 0.160 0.030 0.096 0.116 0.012 0.060 0.106
$h=4$ 0.020 0.060 0.084 0.024 0.052 0.084 0.028 0.056 0.072
$h=8$ 0.008 0.032 0.048 0.020 0.032 0.048 0.024 0.040 0.072
7.1 $h=2$ 0.044 0.096 0.152 0.044 0.096 0.120 0.014 0.058 0.116
$h=4$ 0.040 0.080 0.108 0.032 0.104 0.116 0.012 0.040 0.088
$h=8$ 0.024 0.044 0.064 0.008 0.016 0.028 0.032 0.060 0.088
7.1 $h=2$ 0.036 0.100 0.112 0.024 0.076 0.104 0.048 0.092 0.132
$h=4$ 0.016 0.052 0.092 0.028 0.048 0.068 0.028 0.068 0.104
$h=8$ 0.012 0.036 0.048 0.032 0.056 0.068 0.004 0.028 0.052
$r=4$ 7.1 $h=2$ 0.032 0.116 0.156 0.040 0.084 0.108 0.018 0.050 0.102
$h=4$ 0.032 0.060 0.088 0.032 0.076 0.088 0.016 0.028 0.080
$h=8$ 0.012 0.032 0.048 0.012 0.032 0.048 0.004 0.032 0.098
7.1 $h=2$ 0.032 0.104 0.136 0.032 0.076 0.092 0.014 0.062 0.104
$h=4$ 0.040 0.100 0.108 0.052 0.092 0.116 0.008 0.028 0.076
$h=8$ 0.024 0.056 0.064 0.008 0.040 0.056 0.024 0.056 0.110
7.1 $h=2$ 0.060 0.096 0.136 0.028 0.080 0.132 0.014 0.048 0.102
$h=4$ 0.036 0.064 0.088 0.008 0.044 0.056 0.024 0.068 0.112
$h=8$ 0.020 0.044 0.060 0.020 0.056 0.080 0.016 0.048 0.094
Table 17: Empirical sizes for $\mathbb{W}_{4}$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level. Testing $\delta(z)$.
polynomial series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.088 0.128 0.064 0.104 0.156 0.044 0.112 0.152
$\mathscr{T}_{n}^{*}$ 0.008 0.028 0.064 0.004 0.022 0.074 0.012 0.036 0.082
$h=4$ $\mathscr{T}_{n}$ 0.044 0.080 0.116 0.016 0.048 0.080 0.040 0.076 0.096
$\mathscr{T}_{n}^{*}$ 0.000 0.024 0.076 0.000 0.040 0.072 0.012 0.036 0.084
$h=8$ $\mathscr{T}_{n}$ 0.008 0.024 0.040 0.004 0.008 0.012 0.004 0.016 0.032
$\mathscr{T}_{n}^{*}$ 0.020 0.064 0.072 0.008 0.020 0.076 0.024 0.052 0.090
7.1 $h=2$ $\mathscr{T}_{n}$ 0.048 0.124 0.160 0.032 0.060 0.108 0.060 0.136 0.172
$\mathscr{T}_{n}^{*}$ 0.000 0.016 0.024 0.004 0.012 0.016 0.016 0.028 0.060
$h=4$ $\mathscr{T}_{n}$ 0.060 0.112 0.132 0.028 0.068 0.104 0.044 0.096 0.128
$\mathscr{T}_{n}^{*}$ 0.000 0.012 0.036 0.000 0.000 0.020 0.008 0.012 0.024
$h=8$ $\mathscr{T}_{n}$ 0.016 0.044 0.056 0.008 0.032 0.056 0.012 0.040 0.068
$\mathscr{T}_{n}^{*}$ 0.000 0.040 0.076 0.004 0.038 0.072 0.006 0.044 0.072
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.104 0.140 0.052 0.116 0.168 0.028 0.104 0.156
$\mathscr{T}_{n}^{*}$ 0.000 0.016 0.068 0.000 0.060 0.076 0.000 0.068 0.094
$h=4$ $\mathscr{T}_{n}$ 0.004 0.016 0.024 0.000 0.004 0.012 0.044 0.068 0.108
$\mathscr{T}_{n}^{*}$ 0.024 0.072 0.100 0.016 0.044 0.080 0.012 0.040 0.088
$h=8$ $\mathscr{T}_{n}$ 0.000 0.028 0.048 0.012 0.032 0.068 0.016 0.036 0.068
$\mathscr{T}_{n}^{*}$ 0.006 0.032 0.084 0.000 0.034 0.078 0.004 0.040 0.076
trigonometric series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.052 0.104 0.144 0.044 0.092 0.116 0.048 0.116 0.160
$\mathscr{T}_{n}^{*}$ 0.004 0.022 0.068 0.016 0.020 0.076 0.004 0.034 0.088
$h=4$ $\mathscr{T}_{n}$ 0.012 0.044 0.076 0.032 0.076 0.116 0.052 0.116 0.132
$\mathscr{T}_{n}^{*}$ 0.004 0.024 0.074 0.004 0.030 0.080 0.004 0.038 0.078
$h=8$ $\mathscr{T}_{n}$ 0.000 0.004 0.004 0.000 0.004 0.008 0.008 0.020 0.032
$\mathscr{T}_{n}^{*}$ 0.016 0.036 0.078 0.008 0.020 0.080 0.016 0.028 0.082
7.1 $h=2$ $\mathscr{T}_{n}$ 0.052 0.108 0.132 0.032 0.084 0.108 0.040 0.112 0.148
$\mathscr{T}_{n}^{*}$ 0.000 0.008 0.024 0.000 0.000 0.012 0.008 0.024 0.056
$h=4$ $\mathscr{T}_{n}$ 0.048 0.076 0.116 0.044 0.072 0.104 0.036 0.088 0.128
$\mathscr{T}_{n}^{*}$ 0.008 0.012 0.028 0.000 0.004 0.012 0.000 0.016 0.044
$h=8$ $\mathscr{T}_{n}$ 0.012 0.060 0.072 0.012 0.024 0.040 0.016 0.040 0.116
$\mathscr{T}_{n}^{*}$ 0.000 0.040 0.072 0.004 0.038 0.068 0.008 0.040 0.092
7.1 $h=2$ $\mathscr{T}_{n}$ 0.056 0.128 0.168 0.044 0.084 0.128 0.040 0.108 0.152
$\mathscr{T}_{n}^{*}$ 0.004 0.030 0.064 0.000 0.044 0.088 0.000 0.046 0.094
$h=4$ $\mathscr{T}_{n}$ 0.016 0.024 0.040 0.008 0.056 0.080 0.044 0.104 0.136
$\mathscr{T}_{n}^{*}$ 0.044 0.088 0.100 0.012 0.030 0.068 0.008 0.040 0.094
$h=8$ $\mathscr{T}_{n}$ 0.032 0.044 0.060 0.008 0.036 0.052 0.020 0.040 0.068
$\mathscr{T}_{n}^{*}$ 0.012 0.036 0.082 0.004 0.028 0.062 0.006 0.032 0.068
Table 18: Empirical sizes for $\mathbb{W}_{4}$, rejection probabilities $\mathscr{T}_{n}$ and $\mathscr{T}_{n}^{*}$ at 1%, 5% and 10% level. Testing $\lambda(z)$.
polynomial series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.056 0.136 0.168 0.048 0.096 0.140 0.060 0.124 0.156
$\mathscr{T}_{n}^{*}$ 0.024 0.048 0.088 0.020 0.048 0.084 0.004 0.036 0.082
$h=4$ $\mathscr{T}_{n}$ 0.040 0.068 0.096 0.040 0.080 0.100 0.032 0.088 0.132
$\mathscr{T}_{n}^{*}$ 0.016 0.032 0.094 0.012 0.032 0.076 0.004 0.052 0.092
$h=8$ $\mathscr{T}_{n}$ 0.020 0.056 0.100 0.024 0.068 0.084 0.040 0.096 0.116
$\mathscr{T}_{n}^{*}$ 0.004 0.036 0.088 0.012 0.036 0.068 0.020 0.036 0.100
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.076 0.108 0.036 0.104 0.136 0.072 0.124 0.168
$\mathscr{T}_{n}^{*}$ 0.016 0.032 0.052 0.012 0.032 0.076 0.020 0.052 0.104
$h=4$ $\mathscr{T}_{n}$ 0.032 0.108 0.152 0.028 0.076 0.120 0.060 0.108 0.136
$\mathscr{T}_{n}^{*}$ 0.016 0.028 0.068 0.008 0.020 0.068 0.020 0.040 0.092
$h=8$ $\mathscr{T}_{n}$ 0.044 0.088 0.100 0.020 0.040 0.106 0.012 0.040 0.096
$\mathscr{T}_{n}^{*}$ 0.020 0.060 0.080 0.012 0.028 0.036 0.012 0.028 0.040
7.1 $h=2$ $\mathscr{T}_{n}$ 0.052 0.136 0.192 0.048 0.116 0.148 0.072 0.184 0.216
$\mathscr{T}_{n}^{*}$ 0.008 0.048 0.080 0.012 0.040 0.068 0.020 0.064 0.108
$h=4$ $\mathscr{T}_{n}$ 0.028 0.064 0.108 0.032 0.076 0.092 0.072 0.144 0.204
$\mathscr{T}_{n}^{*}$ 0.008 0.032 0.072 0.008 0.028 0.056 0.024 0.056 0.128
$h=8$ $\mathscr{T}_{n}$ 0.036 0.072 0.076 0.020 0.076 0.112 0.028 0.064 0.080
$\mathscr{T}_{n}^{*}$ 0.012 0.044 0.088 0.004 0.036 0.076 0.016 0.040 0.086
trigonometric series $n=100$ $n=300$ $n=500$
1% 5% 10% 1% 5% 10% 1% 5% 10%
7.1 $h=2$ $\mathscr{T}_{n}$ 0.040 0.112 0.152 0.048 0.100 0.124 0.044 0.108 0.132
$\mathscr{T}_{n}^{*}$ 0.012 0.036 0.088 0.016 0.040 0.094 0.008 0.058 0.082
$h=4$ $\mathscr{T}_{n}$ 0.048 0.096 0.108 0.012 0.072 0.108 0.056 0.116 0.136
$\mathscr{T}_{n}^{*}$ 0.016 0.044 0.088 0.004 0.038 0.082 0.012 0.048 0.088
$h=8$ $\mathscr{T}_{n}$ 0.032 0.056 0.080 0.028 0.076 0.096 0.028 0.064 0.072
$\mathscr{T}_{n}^{*}$ 0.024 0.036 0.072 0.008 0.036 0.080 0.024 0.036 0.088
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.092 0.136 0.064 0.112 0.172 0.056 0.152 0.200
$\mathscr{T}_{n}^{*}$ 0.008 0.044 0.084 0.004 0.028 0.080 0.016 0.056 0.116
$h=4$ $\mathscr{T}_{n}$ 0.036 0.076 0.104 0.020 0.060 0.084 0.052 0.096 0.112
$\mathscr{T}_{n}^{*}$ 0.016 0.028 0.052 0.012 0.024 0.056 0.008 0.044 0.080
$h=8$ $\mathscr{T}_{n}$ 0.028 0.064 0.092 0.028 0.056 0.092 0.016 0.044 0.094
$\mathscr{T}_{n}^{*}$ 0.004 0.036 0.064 0.012 0.040 0.064 0.008 0.020 0.048
7.1 $h=2$ $\mathscr{T}_{n}$ 0.036 0.128 0.164 0.064 0.116 0.164 0.040 0.112 0.132
$\mathscr{T}_{n}^{*}$ 0.004 0.028 0.076 0.016 0.040 0.072 0.024 0.048 0.092
$h=4$ $\mathscr{T}_{n}$ 0.020 0.076 0.092 0.036 0.088 0.116 0.028 0.088 0.116
$\mathscr{T}_{n}^{*}$ 0.008 0.028 0.048 0.012 0.040 0.084 0.000 0.036 0.088
$h=8$ $\mathscr{T}_{n}$ 0.024 0.052 0.076 0.012 0.044 0.072 0.024 0.048 0.060
$\mathscr{T}_{n}^{*}$ 0.012 0.048 0.090 0.008 0.050 0.086 0.000 0.020 0.072